### v1

# Do MLLMs Really Understand the Charts?

Abstract
Although Multimodal Large Language Models (MLLMs) have demonstrated increasingly impressive performance in chart understanding, most of them exhibit alarming hallucinations and significant performance degradation when handling non-annotated charts [^1] The non-annotated charts are those that require viewers to estimate values using the vertical/horizontal axis scale.. Therefore, a question arises: Do MLLMs really understand the charts? Since a human is capable of understanding charts and estimating the values by visual reasoning, we first carefully establish a comprehensive Chart Reasoning Benchmark (CRBench) to rigorously evaluate the visual reasoning abilities of MLLMs on non-annotated charts. We argue that MLLMs are primarily relying on recognition rather than reasoning to interpret the charts. To steer MLLMs to reasonable chart understanding, we propose ChartReasoner that mimics human behavior by grounding their estimation in chart understanding. Extensive results on the proposed CRBench show that ChartReasnoner-3B/7B achieves superior performance in chart reasoning, even compared to GPT-4o and Gemini-2.5-Flash. More importantly, ChartReasnoner also demonstrates the visual reasoning abilities in general chart comprehension on public benchmarks, leading to significant performance gains and enabling MLLMs to rationally understand the charts. The code and dataset will be publicly available upon publication.
Introduction
Multimodal Large Language Models (MLLMs) (Bai et al. 2025; Comanici et al. 2025; OpenAI et al. 2024a; Lu et al. 2024) now play a pivotal role in the field of Artificial Intelligence, particularly for understanding complex visual data. These models have demonstrated a remarkable ability to process charts, analyze their content, provide insightful explanations, and achieve competitive performance on many existing chart benchmarks (Wang et al. 2024; Masry et al. 2022; Xu et al. 2024b; Masry et al. 2025a; Xia et al. 2025).
However, we discover an intriguing phenomenon: When specific numerical annotations are missing in the charts, in other words, when handling non-annotated charts, the MLLMs’ propensity for hallucination increases dramatically—-a failure exemplified in Figure 1, where a model defaults to hallucination in non-annotated charts while our proposed method succeeds through multi-step reasoning. This leads us to a fundamental question: Do MLLMs really understand the charts?
Figure 1: illustration of the visual reasoning deficit in MLLMs when processing non-annotated charts. A “standard Initial Model”, limited by its underdeveloped visual reasoning capacity, often resorts to guessing and fails. In contrast, our “Optimized Model” (ChartReasoner) executes a deliberate, human-like reasoning chain—identifying the target, locating data based on visual scales, and forming a conclusion—to successfully estimate the value.
To investigate this issue, we elaborately introduce a Chart Reasoning Benchmark (CRBench), a novel benchmark that specifically requires visual reasoning for chart understanding, focused on the task of numerical estimation on non-annotated charts. Our goal was to rigorously evaluate the visual reasoning capabilities of current MLLMs on chart understanding and explore whether they truly comprehend the visual clues in the charts. The evaluation reveals that not only open-source MLLMs (Bai et al. 2025; Zhu et al. 2025; Lu et al. 2024) face performance degradation, but even powerful close-source MLLMs, such as GPT-4o (OpenAI et al. 2024a) and Gemini-2.5-Flash (Comanici et al. 2025), also struggle significantly with CRBench. Consequently, we argue that most MLLMs are much more relying on visual recognition capabilities, e.g., Optical Character Recognition (OCR), to read the charts, rather than employing genuine visual reasoning to truly understand them.
Inspired by the success of Reinforcement Learning (RL) in enhancing textual reasoning in domains like mathematics and coding (DeepSeek-AI et al. 2025; OpenAI et al. 2024b), we introduce ChartReasoner, a new model forged through a novel Reinforcement Finetuning (RFT) strategy. This two-stage framework is designed not just to recognize the charts, but also understand the charts via visual reasoning. It begins with a “cold start” through Chain-of-Thought Supervised Fine-Tuning (CoT-SFT) to activate a structured thinking paradigm within the MLLM. This initial ability is then dramatically amplified in the second stage, where the Group Relative Policy Optimization (GPRO) (Shao et al. 2024) training method selectively reinforces correct reasoning pathways and penalizes flawed logic. The experimental results demonstrate that this strategy significantly enhances the visual reasoning capabilities of ChartReasoner, achieving superior performance on the CRbench. Besides, we discovered that the foundational visual reasoning skills cultivated by our method generalize broadly, leading to substantial performance gains across a diverse suite of public chart understanding benchmarks. CharReasoner exhibits highly potential in moving towards a more genuine form of chart comprehension.
The main contributions are summarized as follows:
  • We introduce CRBench, a distinctive benchmark designed to isolate and evaluate genuine visual reasoning capability in chart understanding. It overcomes the limitations of prior work by focusing exclusively on numerical estimation tasks, thus disentangling reasoning from text recognition.
  • We propose ChartReasoner, a powerful MLLM with significantly enhanced visual reasoning capabilities for chart understanding. It achieves excellent performance on our challenging CRBench benchmark, compared with chart-specific and general MLLMs, even surpassing powerful proprietary models like GPT-4o and Gemini-2.5-Flash.
  • We demonstrate that the visual reasoning ability cultivated by our method is foundational and highly generalizable. ChartReasoner is not confined to the specific numerical estimation task, but achieves substantial performance gains across a diverse suite of public, multi-task chart understanding benchmarks.
Related Work
Chart Understanding
The advent of MLLMs, encompassing both powerful open-source and close-source (Comanici et al. 2025; OpenAI et al. 2024a; Team et al. 2025), has laid a foundational groundwork for automated chart understanding. Driven by large-scale visual instruction tuning (Liu et al. 2023), this progress has spurred the development of numerous chart-specific fine-tuned models (Han et al. 2023; Zhang et al. 2024; Masry et al. 2024, 2025b) tailored to enhance performance on these specialized tasks. However, the predominant training paradigm for these models relies on SFT (Liu et al. 2023; Wei et al. 2022), which, as we argue, can lead to superficial mimicry rather than genuine reasoning. In contrast, our work explores a paradigm based on RL (Ouyang et al. 2022) to cultivate a deeper, more robust form of chart understanding.
Figure 2: The training paradigm of ChartReasoner and the data distribution of CRBench. A: ChartReasoner leverages a two-stage RFT strategy. Stage 1 activates the model’s reasoning abilities via SFT on CoT data, while Stage 2 uses GRPO with a multi-component reward system to reinforce correct chart understanding. B: The composition of CRBench, detailing the distribution of seven chart types across both synthetic and real data sources.
Methods Synthetic Charts Real Charts Overall
Box Area Radar Scatter Bar Line Combo Bar Line Combo
Human 74.72 56.70 83.97 77.90 90.10 93.71 77.27 80.36 65.22 52.00 78.64
Open-source Models
InternVL3-2B (Zhu et al. 2025) 31.87 7.32 5.77 23.19 51.40 45.70 32.98 39.29 33.04 20.80 31.59
Qwen2.5-VL-3B (Bai et al. 2025) 49.45 15.24 19.23 52.74 74.58 82.45 62.48 67.86 54.78 48.80 57.68
Ovis1.6-llama3.2-3B (Lu et al. 2024) 16.48 2.44 10.90 14.22 17.04 12.25 11.52 12.50 13.04 9.60 12.47
Gemma-3-4B (Team et al. 2025) 15.38 3.05 10.90 10.28 8.38 8.94 6.28 5.36 9.57 8.80 8.32
Qwen2.5-VL-7B (Bai et al. 2025) 72.53 28.05 26.92 66.08 84.08 86.42 71.03 63.39 60.87 47.20 66.25
InternVL3-8B (Zhu et al. 2025) 53.85 19.51 14.10 43.33 62.29 61.92 48.52 58.93 55.65 42.40 42.40
Close-source Models
GPT-4o (OpenAI et al. 2024a) 32.97 15.24 12.82 25.82 22.35 25.17 19.90 15.18 26.96 16.80 21.69
Gemini-2.5-Flash (Comanici et al. 2025) 72.53 32.93 11.54 59.96 70.11 71.19 55.50 53.57 46.96 37.60 55.32
Chart-Specific Models
ChartGemma-3B (Masry et al. 2025b) 10.99 10.98 7.05 21.44 42.74 32.78 24.43 37.50 42.61 28.80 26.74
TinyChart-3B (Zhang et al. 2024) 13.19 7.93 7.69 25.16 56.15 54.30 36.65 57.14 40.87 34.40 35.83
ChartInstruct-7B (Masry et al. 2024)
10.99 1.22 4.49 16.63 35.47 16.56 18.85 51.79 45.22 18.40 20.91
ChartVLM-7.3B (Xia et al. 2025) 9.89 10.37 7.69 10.28 70.39 45.36 32.81 50.00 54.78 35.20 33.63
ChartLlama-13B (Han et al. 2023) 10.99 1.83 5.77 5.25 3.35 4.30 4.01 8.04 9.57 7.20 5.01
Table 1: Performance evaluation on CRBench with direct answering, covering a range of open-source, close-source, and chart-specific models. This table shows the results when models are prompted to directly output the final answer without intermediate reasoning steps.
A suite of chart benchmarks, including ChartQA (Masry et al. 2022), PlotQA (Methani et al. 2020), and others (Xu et al. 2024b; Han et al. 2023; Wang et al. 2024), have been proposed to evaluate MLLMs on diverse chart tasks. While instrumental, their task design often prioritizes language-level reasoning (Lewkowycz et al. 2022) (e.g., calculating averages or summarizing) over fundamental visual reasoning (Johnson et al. 2017; Hudson and Manning 2019). The core skill of grounding reasoning in a chart’s geometry, such as estimating values from axes on non-annotated charts, remains largely untested. Our CRBench is specifically designed to isolate and evaluate this crucial visual reasoning capability.
Visual Reasoning
RL has been successfully employed to significantly enhance the reasoning abilities of Large Language Models (LLMs) (DeepSeek-AI et al. 2025), particularly in complex domains like mathematics and coding (Shao et al. 2024). Inspired by this success in language-only models, several works have begun to apply similar RL-based paradigms to MLLMs (Feng et al. 2025; Tan et al. 2025; Huang et al. 2025), successfully activating their visual reasoning capabilities on various challenging tasks (e.g., visual counting, spatial transformation). Building on these advancements, our work adapts and applies an RL-based framework to the specialized domain of chart understanding.
Chart Reasoning Benchmark
We introduce Chart Reasoning Benchmark (CRBench), a comprehensive, multi-domain, and reasoning-centric benchmark designed to rigorously assess the visual interpretation capabilities of MLLMs on charts that lack explicit numerical annotations. Engineered to move beyond simple OCR-dependent tasks, the benchmark comprises a total of 2,453 question-answer pairs. It features a majority (2,101 pairs) of synthetically generated charts to ensure controlled complexity and a significant portion (352 pairs) sourced from real-world examples to guarantee practical relevance. The benchmark provides extensive coverage across seven primary chart types, including bar, line, scatter, and combo charts, with a detailed breakdown of the data distribution shown in Figure 2. This structural diversity is complemented by thematic breadth, with data spanning 38 distinct topics, including finance, healthcare, and technology, etc. This dual emphasis on structural and thematic variety ensures a rigorous evaluation, mitigating the risk of models overfitting to specific chart formats or familiar domains.
Data Curation
Our data generation process integrates automated synthesis with manual curation to ensure both scale and quality.
Synthetic Chart Generation.
Our synthetic chart generation process is partially adapted from the Code-as-Intermediary Translation (CIT) methodology proposed by (He et al. 2024), where executable plotting code serves as the ground truth for each chart. The process begins with a curated set of seed scripts, which are then programmatically diversified using Self-Instruct (Wang et al. 2023) and Evol-Instruct (Xu et al. 2024a) techniques to generate a vast library of visually complex charts. A critical constraint is the deliberate omission of numerical labels on data points, ensuring that every chart necessitates visual estimation. To maximize yield, a self-repair mechanism leverages an LLM to debug and correct any code that fails during execution. Following an automated visual fidelity check by a multimodal language model, the entire collection of synthesized data underwent a final, rigorous human review. This manual verification step served to confirm the high quality of the chart images and the accuracy of their corresponding question-answer pairs. This code-centric approach, combined with multiple stages of validation, provides an unimpeachable ground truth, allowing us to generate verifiably correct Q&A pairs.
Real Chart Collection.
To anchor our benchmark in real-world applications, we sourced charts from reputable data repositories such as Statista and Our World in Data. Each chart was manually vetted by human annotators to meet strict criteria: high visual quality, data integrity, and a complete absence of explicit numerical annotations. Following selection, an MLLM was used to generate candidate question-answer pairs for each chart. Every MLLM-generated pair then underwent a final round of human verification and refinement to guarantee the accuracy and relevance of both the question and its ground-truth answer.
Evaluation Protocol
Standard exact-match accuracy is ill-suited for value estimation from non-annotated charts, as it fails to account for the slight perceptual ambiguity inherent in the task, even for human observers. To address this, we employ a relaxed accuracy metric, which judges a prediction correct if its relative error from the ground-truth value falls within a tolerance threshold, denoted as $\tau$. To align this threshold with human performance, we conducted an empirical study and found that human estimations consistently fall within a 2% error margin. Accordingly, we empirically set $\tau=0.02$. Formally, a model’s predicted value, $A_{pred}$, is deemed correct if and only if it satisfies the following condition relative to the ground truth, $A_{gt}$:
$A_{pred}\in[(1-\tau)\times A_{gt},(1+\tau)\times A_{gt}]$
This protocol ensures that our evaluation is both rigorous and fairly aligned with human-level interpretive capabilities, rewarding models for precise visual reasoning rather than penalizing them for minor, human-like estimation variance.
Methods Synthetic Charts Real Charts Overall
Box Area Radar Scatter Bar Line Combo Bar Line Combo
Open-source Models
InternVL3-2B (Zhu et al. 2025) 25.27 8.54 9.62 25.16 51.68 43.05 34.55 43.75 34.78 32.00 32.98
Qwen2.5-vl-3B (Bai et al. 2025) 46.15 14.63 17.95 52.30 72.35 80.79 62.65 66.96 53.04 46.40 56.62
Ovis1.6-llama3.2-3B (Lu et al. 2024) 12.09 3.66 7.69 14.66 13.13 10.93 11.69 8.04 16.52 12.00 11.66
Gemma-3-4B (Team et al. 2025) 9.89 3.05 12.82 11.82 9.22 12.25 6.81 8.04 12.17 7.20 9.34
Qwen2.5-vl-7B (Bai et al. 2025) 69.23 20.12 19.87 61.93 74.02 83.77 67.71 49.11 53.04 46.40 60.74
InternVL3-8B (Zhu et al. 2025) 36.73 12.80 12.18 39.17 43.58 47.68 36.82 38.39 46.09 34.40 36.73
Close-source Models
GPT-4o (OpenAI et al. 2024a) 28.57 12.20 11.54 25.61 21.23 27.15 18.15 13.39 26.96 18.40 20.87
Gemini-2.5-Flash (Comanici et al. 2025) 68.13 25.61 7.69 61.93 72.07 75.17 55.85 49.11 52.17 39.20 55.77
Chart-Specific Models
ChartReasoner-3B (Ours) 58.24 21.95 22.44 56.24 79.61 86.42 66.14 65.18 54.78 48.80 61.27
ChartReasoner-7B (Ours) 72.53 31.71 32.05 65.43 82.96 84.44 71.55 65.18 66.09 53.60 67.06
Table 2: Comparison of ChartReasoner with representative MLLMs on the porposed CRBench, using Chain-of-Thought (CoT) prompting to elicit the models’ intermediate reasoning steps. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
ChartReasoner
We propose ChartReasoner, a MLLM designed to perform visual reasoning for better visual understanding on non-annotated charts. We formally define this task as follows: given a chart image $I$, and a corresponding textual question $Q$, the goal is to derive a numerical answer $A$ with a reasoning procedure $R$. This process can be represented as a mapping function $\mathcal{F}$:
$\mathcal{F}:(I,Q)\rightarrow(R,A)$
where $I$ is the chart image, $Q$ is the question in text, $R$ is the step-wise reasoning procedure in text, and $A\in\mathbb{R}$ is the numerical answer. The fundamental challenge lies in interpreting non-annotated charts, which requires the model to reason about geometric structures (e.g., axes, scales, positions) to infer values, rather than simply extracting them via text recognition. While the mapping $\mathcal{F}$ inherently performs this reasoning, producing an explicit reasoning chain $R$ that explains the derivation is an optional, secondary objective.
To address this challenge, we propose a two-stage Reinforcement Finetuning (RFT) strategy. This approach is designed to first instill a robust, human-like reasoning framework and then meticulously refine the model’s numerical precision. As illustrated in Figure 2, the RFT pipeline consists of two sequential stages: (1) Reasoning Activation via SFT, which uses supervised fine-tuning to teach the model the structure of reasoning, followed by (2) Reasoning Generalization via RL, which uses reinforcement learning to improve the accuracy and generalizability.
Stage 1: Reasoning Activation via SFT
The initial stage of our pipeline aims to establish a foundational reasoning paradigm. Instead of having the model directly guess an answer, we teach it to adopt a structured, step-by-step thought process that mirrors human analysis. To achieve this, we first generate a high-quality dataset by distilling detailed Chain-of-Thought (CoT) processes from a MLLM with emerging reasoning abilities. Fine-tuning our base MLLM on these 68,500 synthesized samples teaches the model to emulate a human’s analytical approach, guiding it to identify and utilize critical chart components—such as axes, scales, and legends—before arriving at a final answer.
Formally, we employ SFT on this dataset. Each data instance is a tuple $(x,q,r,a)$, where $x$ is the chart image, $q$ is the question, $r$ is the intermediate reasoning chain, and $a$ is the final answer. The training objective is to maximize the log-likelihood of the model generating the complete sequence $y$ (the concatenation of $r$ and $a$) given the image $x$ and question $q$:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,q,r,a)\sim\mathcal{D}}\sum_{t=1}^{|y|}\log\pi_{\theta}(y_{t}|x,q,y_{<t})$ (1)
where $\mathcal{D}$ is our CoT dataset and $\pi_{\theta}$ is the policy of the model with parameters $\theta$. The resulting fine-tuned model, denoted as $\pi_{\text{SFT}}$, learns a robust template for visual reasoning and serves as the starting point for the next stage.
Stage 2: Reasoning Generalization via RL
Building on the reasoning foundation from Stage 1, the second stage focuses on enhancing the model’s precision and reliability for the numerical estimation task. For this, we use a smaller, high-signal dataset of 3,400 samples curated to target the model’s specific weaknesses. These samples are identified by selecting problems where the SFT-tuned model exhibits ”stochastic correctness”—that is, problems it can solve but not consistently. By training on these borderline cases with higher-resolution images, we force the model to refine its visual interpretation skills.
We employ GRPO (Shao et al. 2024), an efficient and scalable reinforcement learning algorithm, to fine-tune the policy $\pi_{\text{SFT}}$. Unlike traditional algorithms like PPO (Schulman et al. 2017), GRPO forgoes a computationally expensive value network and instead calculates relative advantages by comparing rewards within a group of sampled responses. For each input $(x,q)$, we sample a group of $G$ candidate answers $\{a_{1},a_{2},\dots,a_{G}\}$ from the current policy $\pi_{\beta}$. Each answer $a_{i}$ receives a reward $R(a_{i})$, and these rewards are used to compute a normalized relative advantage $A_{i}$ for each sample:
$A_{i}=\frac{r_{i}-\text{mean}\{r_{1},\dots,r_{G}\}}{\text{std}\{r_{1},\dots,r_{G}\}}$ (2)
The policy is then updated to increase the probability of actions with positive advantages, while a KL-divergence penalty against the reference model $\pi_{\text{SFT}}$ ensures stable training.
Reward Function Design
The effectiveness of our RL stage hinges on a well-designed reward function. Our function $R(a_{i})$ is a composite of two components, targeting both response structure and numerical accuracy:
$R(a_{i})=R_{\text{format}}(a_{i})+R_{\text{acc}}(a_{i})$ (3)
Format Reward.
To encourage interpretable and well-structured outputs, we provide a binary format reward, $R_{\text{format}}$. The model receives a reward of 1 if its response strictly adheres to our predefined template—reasoning enclosed in <think></think> tags and the final answer in <answer></answer> tags—and 0 otherwise.
Accuracy Reward.
To overcome the sparse signal from a simple correct/incorrect binary reward, we introduce a continuous accuracy reward, $R_{\text{acc}}$. This reward provides a fine-grained signal that recognizes ”nearly correct” answers. For a predicted answer $A_{\text{pred}}$ and a non-zero ground truth $A_{\text{gt}}$, we first calculate the relative error:
$d_{\text{rel}}=\frac{\lvert A_{\text{pred}}-A_{\text{gt}}\rvert}{\lvert A_{\text{gt}}\rvert}$ (4)
Based on a human-calibrated tolerance threshold of $\varepsilon=0.02$, we define the reward using a piecewise quadratic function that smoothly decays from 1 to 0:
$R_{\text{acc}}(d_{\text{rel}})=\begin{cases}\left(1-\frac{d_{\text{rel}}}{\varepsilon}\right)^{2}&\text{if }d_{\text{rel}}<\varepsilon\\ 0&\text{if }d_{\text{rel}}\geq\varepsilon\end{cases}$ (5)
We employ the quadratic formulation for two critical reasons. First, this design provides a clear, bounded, and intuitive reward range. It yields a reward of 1 for a perfect answer ($d_{\text{rel}}=0$) and smoothly decay to 0 as the relative error hits the 2% tolerance boundary. Second, the quadratic shape creates a desirable non-linear decay. It has a gentle slope for subtile errors, granting substantial partial credit for close answers, while the penalty accelerates as the error approaches the tolerance threshold. This behavior encourages the model to make fine-grained improvements when it is already close to the correct answer, while strongly penalizing larger, unacceptable deviations.
Figure 3: Qualitative analysis on CRBench. GPT-4o and Qwen2.5-VL-7B exhibits hallucination in the answer and reasoning procedure, respectively. In contrast, our ChartReasoner-7B is able to produce a coherent and correct step-by-step reasoning process, leading to an accurate answer.
Methods CRBench CharXiv (R) CharXiv (D) ChartBench ChartQAPro
ChartGemma-3B (Masry et al. 2025b) 26.74 12.50 21.30 - 6.84
TinyChart-3B (Zhang et al. 2024) 35.83 8.30 16.15 - 13.25
ChartInstruct-7B (Masry et al. 2024) 20.91 8.80 21.40 - 4.88
ChartVLM-7.3B (Xia et al. 2025) 33.63 - - 12.06 -
ChartLlama-13B (Han et al. 2023) 5.01 14.20 19.23 21.30 -
ChartReasoner-7B (Ours) 67.06 43.40 63.90 45.34 36.06
Table 3: Performance of ChartReasoner compared to other chart-specific models on various public chart understanding benchmarks. The results of the compared methods are from the corresponding benchmarks. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
Experiments
Experimental Setups
Training Setting
The implementation was built upon the ModelScope SWIFT framework (Zhao et al. 2025) for all training procedures. We initialize ChartReasoner with Qwen2.5-VL (Bai et al. 2025) and tune it with the dataset described in the previous section. All experiments were carried out on a server with 8 NVIDIA H800 GPUs. The methods evaluated on CRBench follow their default inference settings. All results were obtained from a single forward pass with a fixed random seed of 42. Please refer to the supplementary materials for more details.
Training Paradigms and Baselines
To systematically evaluate our training strategies, we conduct a comparative study in four distinct paradigms.
  • CoT-SFT: Provides a reasoning “cold start” by fine-tuning the model on CoT dataset, with the reasoning data distilled from Qwen2.5-VL-32B (Bai et al. 2025).
  • GRPO-Zero: An ablation study in which our GRPO algorithm is applied directly to the base model to assess the efficacy of reinforcement learning in isolation.
  • RFT: Our proposed final paradigm starts from the CoT-SFT checkpoint and uses GRPO to further refine and generalize the model’s reasoning capabilities.
To further validate the generalization and effectiveness of our training paradigms, we conduct experiments across a diverse set of base models. Specifically, we evaluate our methods on both the Qwen2.5-VL-Instruct (Bai et al. 2025) series, comparing the 3B and 7B variants to assess scalability.
Methods Accuracy
Synthetic Real Overall
ChartReasoner-3B
Zero-Shot 46.41 42.90 45.90
+CoT-SFT 51.02 45.45 50.22
+GRPO-Zero 56.59 56.82 56.62
+RFT 60.92 54.83 60.05
ChartReasoner-7B
Zero-Shot 60.54 45.17 58.34
+CoT-SFT 63.59 51.70 61.88
+GRPO-Zero 66.44 54.83 64.78
+RFT 66.97 60.23 66.00
Table 4: Ablation study showing the individual contribution of each strategy (CoT-SFT, GRPO-Zero, and RFT) to the final performance of the ChartReasoner models. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
Experimental Results
Evaluation on CRBench
The results in Table 1 underscore the significant challenge our CRBench presents to current models. The generally low scores across the board reveal a critical weakness: the widespread inability to reason over non-annotated charts. This difficulty stems from our benchmark’s emphasis on genuine visual reasoning—inferring values from axes and scales—rather than relying on annotation-based shortcuts common in other benchmarks. Furthermore, while most methods obtains comparable results on synthetic and real charts, the overall scores are still unsatisfying. To reveal the failure stems, we employed CoT prompting to compel models to externalize their reasoning process. The results in Table 1 and Table 2 show that most methods do not benefit from CoT prompting. It can be attributed to the lack of visual reasoning ability, as shown in Figure 3.
Evaluation of ChartReasoner
Our comprehensive evaluation confirms that ChartReasoner-7B decisively outperforms not only existing chart-specific models but also top-tier proprietary models like GPT-4o and Gemini-2.5-Flash, as shown in Table 2. This success stems from its ability to perform genuine visual reasoning, a skill that other models lack, as evidenced by their reliance on taking annotation shortcuts. This advanced capability is particularly evident in its handling of complex chart typologies. For instance, on challenging radar charts, where many models falter, ChartReasoner demonstrates a significant performance leap, proving it can interpret complex visual structures.
Ablation Analysis of the Training Paradigm
Our experimental results, summarized in Table 4, strongly validate the effectiveness of the RFT strategy for enhancing visual reasoning. As discussed in the previous section, simply prompting a model with a CoT format often leads to performance degradation, as it forces the model to articulate a reasoning process it does not possess. Our training paradigms were designed to address this fundamental issue. The result clearly indicates that while a naive CoT-SFT offers only a marginal improvement, it is the synergistic combination within the RFT model that yields the most significant performance gains, consistently outperforming all other configurations.
Case Study of Visual Reasoning
In Figure 3, powerful models like GPT-4o and Qwen2.5-VL-7B either misinterpret the query or resort to factual hallucination. In contrast, our ChartReasoner, sculpted by the RFT strategy, demonstrates a flawless, step-by-step reasoning process. This case study provides a compelling visual proof of our quantitative findings: ChartReasoner enhanced by the RFT strategy is transformed from a system prone to errors and hallucination into a reliable and structured visual reasoner that moves beyond merely optimizing for the answer.
Evaluation of ChartReasoner on the Public Benchmarks
The robust generalization of ChartReasoner is confirmed by its strong performance across a suite of public benchmarks, as detailed in Table 3. This success stems from our model’s core visual reasoning capability, honed through our training strategy, which is highly transferable to other reasoning-intensive tasks. This approach contrasts with that of other chart-specific models relying on SFT. Although SFT can improve performance on specific tasks, it often fails to instill a generalizable reasoning faculty, limiting extensibility. In contrast, ChartReasoner, enhanced by the RFT strategy, is specifically designed to cultivate this fundamental reasoning, validating the superiority of our method.
Conclusion
In this paper, we investigate a compelling yet significant question: ”Do MLLMs really understand the charts?” By establishing the CRBench, we extensively evaluated open-source, close-source, and chart-specific MLLMs. The results shows a significant degradation in the performance of these models, and, through chain-of-thought reasoning, revealed their inability to estimate numerical values through visual reasoning, similar to human behavior. To address this issue, we propose ChartReasoner, which enhances the visual reasoning ability of MLLMs via an RFT strategy. This strategy first activates reasoning capabilities through SFT, and then generalizes reasoning abilities through RL. Extensive experiments conducted on the proposed CRBench and public chart reasoning datasets demonstrate the effectiveness of ChartReasoner. This work paves the way for empowering MLLMs to really understand the charts in a human-like manner.
Appendix
  • Sec. A provides details of the proposed benchmark, CRBench, including the task definition, the data construction pipeline, and data sources, topics, and chart types.
  • Sec. B describes the training strategy of the ChartReasoner, detailing the construction of the SFT and RL datasets.
  • Sec. C presents further details on the inference.
  • Sec. D includes other relevant details and extended results to supplement the main findings.
Appendix A A. CRBench Benchmark Details
Problem Definition
The primary task addressed by our benchmark, CRBench, is numerical value estimation on non-annotated charts. Formally, given a chart image $C$ and a query $Q$ that specifies a target data point within the chart, the goal is to produce a numerical answer $A$ that accurately estimates the value of that data point. Crucially, the chart image $C$ is non-annotated, meaning that the numerical values corresponding to graphical elements (e.g., the height of a bar, a point on a line) are not present as explicit text labels.
This task is fundamentally a visual reasoning problem, rather than a simple recognition or textual reasoning task. To arrive at the correct answer $A$, a model cannot only rely on Optical Character Recognition (OCR). Instead, it must perform a multi-step cognitive process grounded in the visual geometry of the chart:
  1. 1.
    Semantic Understanding & Grounding: The model must first parse the query $Q$ and correctly associate the textual description with the corresponding graphical elements in the chart image $C$ (e.g., a specific bar, a specific colored line).
  2. 2.
    Structural and Scale Interpretation: The model must identify and interpret the chart’s structural components, particularly the relevant axes (e.g., the y-axis) and their numerical scales, including the range and the value represented by tick marks and grid lines.
  3. 3.
    Spatial and Proportional Reasoning: Finally, the model must perform spatial reasoning by comparing the target graphical element’s dimension (e.g., its height or position) against the interpreted scale of the axis. This often requires proportional estimation or interpolation between labeled tick marks to infer the final numerical value.
By designing a task that necessitates this entire reasoning chain, we directly evaluate a model’s ability to not just recognize a chart, but to truly understand its underlying quantitative information.
Figure 4: The automated data generation pipeline for CRBench.
Data Generation Pipeline
The data generation pipeline for CRBench and training datasets, visualized in Figure 4, is an automated, multi-stage process designed to programmatically create a large and diverse set of non-annotated charts. The entire workflow is built upon generating and refining executable Python plotting scripts, ensuring that each data sample is verifiably correct.
  1. 1.
    Chart Generation by Large Language Model (LLM): The process begins with a curated seed pool of high-quality plotting scripts that serve as few-shot demonstrations. Using a Self-Instruct methodology, an LLM is then prompted to programmatically generate a vast and diverse set of new, non-annotated chart plotting scripts across 38 distinct topics.
  2. 2.
    Validation and Evolutionary Enhancement: Each generated script is rigorously validated via sandboxed execution to ensure it runs error-free. A key feature of the pipeline is an AI-driven repair loop, where the language model automatically debugs and corrects any code that fails execution. Subsequently, a portion of these validated scripts undergo an optional Evol-Instruct phase to further increase their visual and structural complexity.
  3. 3.
    Final Curation: Only scripts that successfully pass the entire validation and enhancement pipeline are accepted into the final datasets. These validated scripts are then executed to generate the final, high-fidelity chart images used in our benchmark and datasets.
Ground-Truth Q&A Generation and Verification.
After generating the chart images, we create the corresponding ground-truth Q&A pairs. We first use a LLM to generate questions and answers based on the chart’s source code. Each generated Q&A pair then undergoes a rigorous, multi-tiered verification: it is first cross-checked for accuracy by another LLM instance, and subsequently undergoes a final manual review by human annotators to guarantee its unequivocal correctness.
Data topics and Chart Examples
To ensure the breadth relevance of CRBench, the data generation process samples from a diverse set of 38 distinct topics, span a wide range of domains including finance, healthcare, technology, and social sciences, guaranteeing that the charts cover a variety of contexts and narratives. Furthermore, to robustly evaluate a model’s visual reasoning capabilities across different graphical representations, CRBench incorporates seven primary chart types. Figure 6 provides a representative example for each of these types, showcasing the visual diversity and complexity present in our benchmark.
Real-World Chart Collection
To complement our synthetic dataset and ensure CRBench reflects the challenges of real-world applications, we curate a substantial collection of charts from public online sources. This process involves careful sourcing from key data providers, a rigorous manual filtering stage, and a hybrid human-AI approach to generate high-quality question-answer pairs.
Statista
A significant portion of the real-world charts is sourced from Statista [^2] https://www.statista.com/, a leading global platform specializing in market and consumer data. Statista provides professional, data-driven visualizations for a wide array of industries, covering topics from economic indicators and market forecasts to technology trends and consumer behavior. The charts on this platform are typically clean, well-structured, and designed for clear communication in business and academic contexts.
The World in Data
The second major source is Our World in Data [^3] https://ourworldindata.org/, a renowned scientific online publication based at the University of Oxford. Its mission is to make data and research on the world’s most significant challenges, such as global health, economic development, and environmental change, accessible and understandable. The visualizations from the World in Data are often more complex and data-rich, including long-term time-series line charts, multi-country scatter plots, and stacked area charts.
In addition to these two primary repositories, the collection is supplemented by charts from various other miscellaneous public reports and online publications. All data collection is conducted in strict adherence to the copyright policies, terms of service, and licensing agreements of each source to ensure full ethical compliance.
Manual Filtering and Vetting.
A large pool of charts is initially collected. Each chart then undergoes a meticulous manual vetting process conducted by our team of annotators to ensure it met the stringent criteria for inclusion in CRBench. A chart is only accepted if it satisfied all of the following conditions:
  1. 1.
    High Visual Quality: The image has to be of sufficient resolution, clear, and free of significant compression artifacts or other visual noise that could impede interpretation.
  2. 2.
    Data Integrity: The chart has to be coherent and visually consistent, with clearly defined axes, legends, and graphical elements.
  3. 3.
    Absence of Annotations: We exclusively select charts where numerical values are not explicitly printed on the graphical elements (e.g., no numbers on top of bars or next to points on a line). This constraint is fundamental to our benchmark, as it forces a model to perform genuine visual reasoning rather than relying on OCR shortcuts.
Charts that fails to meet any of these standards are discarded from the final dataset.
Question-Answer Pair Generation.
Once a real-world chart is approved, we employ a two-stage, human-in-the-loop process to generate its corresponding question-answer pair:
  1. 1.
    MLLM-based Candidate Generation: We use a capable Multimodal Large Language Model (MLLM) to generate an initial set of candidate question-answer pairs for each chart. The model is prompted to ask a specific numerical estimation question based on the visual information in the chart.
  2. 2.
    Human Verification and Refinement: Every MLLM-generated pair is subsequently subjected to rigorous human review. Annotators would first verify the question’s clarity and relevance. Then, they would carefully perform the visual estimation themselves to validate the accuracy of the answer. If necessary, they would refine the question’s phrasing or correct the ground-truth answer to ensure the final Q&A pair was unambiguous and factually correct.
This careful, multi-stage curation process ensures that the real-world portion of CRBench is not only challenging and relevant but also contains reliable and accurate ground-truth data. Figure 7 showcases several examples of the final curated real-world charts from our collection, along with their corresponding question-answer pairs.
Appendix B B. Details of Datasets and Training
CoT Data Construction for Reasoning Activation
The dataset used for the Supervised Fine-Tuning (SFT) ”cold start” phase is meticulously constructed through a knowledge distillation process. The goal is to generate high-quality Chain-of-Thought (CoT) data that can effectively activate the reasoning paradigm of our base models. This process leverages a powerful teacher model(In our experiment, we use the Qwen2.5-VL-32B-Instruct as the teacher mdoel) to generate reasoning traces for the question-answer pairs in our CRBench dataset.
The data construction pipeline, implemented in the generation scripts, involves several key stages to ensure the quality and validity of the final CoT samples:
  1. 1.
    Prompting for Chain-of-Thought Generation: For each (C,Q,A) triplet from original data, we prompt the teacher model to generate a step-by-step reasoning process. Critically, the prompt(in Figure8) is designed to guide the model to derive the provided ground-truth answer, forcing it to articulate a logical pathway that leads to the correct result. The model is required to structure its output using specific tags, separating the reasoning steps (<think>...</think>) from the final answer (<answer>...</answer>).
  2. 2.
    Validation and Filtering: Each generated CoT sample undergoes a rigorous, multi-step validation process to filter out low-quality or incorrect reasoning:
    • Structural Check: The generated text is first parsed to ensure that both the reasoning and answer tags are present. Samples with missing tags are discarded.
    • Answer Verification: The final answer extracted from the <answer> tag is programmatically compared against the ground-truth answer. We employ a robust evaluation function that checks for both exact string matches and numerical equivalence within a tolerance threshold to ensure correctness.
    • Leakage Detection: The generated reasoning trace within the <think> tags is scanned for any mention of the ”original answer.” This crucial step prevents the model from ”cheating” by simply copying the ground-truth answer into its reasoning, ensuring that the generated thought process is genuine.
This meticulous, multi-stage distillation and verification process—with examples showcased in Figure 9 and Figure 10—ensures that the data used for the reasoning activation phase is both correct and logically sound, providing a strong foundation for the subsequent reinforcement learning stage.
Dataset Construction for GRPO
The dataset for the Generative Rejection-based Policy Optimization (GRPO) stage is meticulously curated to focus the training on the most informative examples. Instead of using a random subsample of data, we construct a specialized dataset of moderately difficult samples. These are question-answer pairs that the base model can sometimes answer correctly but frequently gets wrong, representing the frontier of its reasoning capabilities and thus offering the richest signal for reinforcement learning.
The curation process inspired by the principles of rejection sampling and active learning, involves a multi-round, varied-prompting inference pipeline designed to probe the model’s knowledge boundaries. We first run inference on the entire CRBench dataset multiple times using the same base model. To elicit diverse responses, we systematically vary the prompts between rounds (e.g., alternating between direct answering and Chain-of-Thought prompts) and increase the model’s sampling temperature to 0.9. The correctness of every generated response is logged. After the evaluation is complete, we filter this log to isolate the key samples: we select only those question-answer pairs that the model answered correctly in at least one round but incorrectly in at least one other round.
The rationale behind this selective filtering strategy is to maximize the efficiency and impact of the subsequent RL training. By focusing on these ”boundary” cases, we avoid wasting computational resources on problems the model has already mastered (always correct) or those that are currently far beyond its reach (always incorrect). Training on these ”boundary” samples forces the policy to learn to distinguish between successful and flawed reasoning pathways for the exact same problem. This targeted approach ensures that the GRPO stage is dedicated to resolving ambiguity and reinforcing robust reasoning where it is most needed, leading to more significant and generalizable improvements in the model’s core reasoning abilities.
Detailed Formulation of the Accuracy Reward
A core component of the GRPO framework is the Accuracy Reward ($R_{\text{acc}}$), which is designed to provide a dense, fine-grained signal for optimizing the model’s numerical estimation capabilities. A simple binary reward (correct/incorrect) is often too sparse for reinforcement learning, as it fails to differentiate between a near-miss and a completely wrong answer. To overcome this, we designed a continuous reward function that recognizes and rewards ”nearly correct” answers, thereby creating a smoother optimization landscape. The formulation of this reward is visualized in Figure 5.
Figure 5: Visualization of our quadratic accuracy reward function ($R_{\text{acc}}$) with a tolerance threshold of $\varepsilon=0.02$. The reward smoothly decays from 1 to 0, with a gentle slope for small errors and an accelerating penalty for larger deviations.
For a predicted numerical answer $A_{\text{pred}}$ and a non-zero ground-truth answer $A_{\text{gt}}$, we first calculate the relative error:
$d_{\text{rel}}=\frac{\lvert A_{\text{pred}}-A_{\text{gt}}\rvert}{\lvert A_{\text{gt}}\rvert}$ (6)
Based on a tolerance threshold of $\varepsilon=0.02$, we define the reward using a piecewise quadratic function that smoothly decays from a maximum of 1 to 0:
$R_{\text{acc}}(d_{\text{rel}})=\begin{cases}\left(1-\frac{d_{\text{rel}}}{\varepsilon}\right)^{2}&\text{if }d_{\text{rel}}<\varepsilon\\ 0&\text{if }d_{\text{rel}}\geq\varepsilon\end{cases}$ (7)
Our accuracy reward function provides a dense, informative, and well-behaved signal that is ideally suited for guiding our reinforcement learning process towards generating highly accurate and reliable numerical estimations.
Appendix C C. Inference Details
This section provides additional details regarding our inference and evaluation setup to ensure full reproducibility.
Inference Hyperparameters and Framework
All experiments reported in this paper were conducted using the default hyperparameters of each respective model, with no model-specific tuning performed at inference time. This ensures a fair and consistent comparison across all models. To guarantee reproducibility, the random seed for all generation processes was fixed to 42. Our inference pipeline is built upon the vLLM framework, which provides efficient, high-throughput serving for Large Language Models.
Prompting Strategies for Evaluation
To thoroughly investigate the models’ reasoning capabilities and behaviors, we employed three distinct prompting strategies during evaluation, designed to control the externalization of the model’s thought process. The first is Direct Answering, a baseline prompt that instructs the model to provide only the final answer. The second, Optional Chain-of-Thought (Optional-CoT), encourages but does not require the model to generate a reasoning trace, allowing us to observe its natural tendency to reason. The third, Forced Chain-of-Thought (Forced-CoT), strictly requires a step-by-step reasoning process before the final answer, compelling the model to articulate its logic. As illustrated in Figure 14 and Figure 15, which showcases two comparative examples, these varied strategies allow us to gain deeper insights into not just what the models answer, but how their performance is affected by the explicit demand for reasoning.
Appendix D D. Evaluation Protocols for Specialized Models and Benchmarks
This section outlines the precise met hodologies used to evaluate specialized, chart-specific models and to benchmark our ChartReasoner on existing public datasets. Our protocols were designed to ensure maximum fairness and comparability by strictly adhering to the author-recommended settings and official evaluation scripts for each model and benchmark.
Evaluation of Chart-Specific Models on CRBench
Evaluating the performance of existing chart-specific models on our proposed CRBench required a careful approach to ensure a fair comparison. These models, including ChartGemma, TinyChart, ChartInstruct, ChartVLM, and ChartLlama, are often fine-tuned on specific data formats and are highly optimized for a particular style of prompting.
For this reason, we did not use our generalized prompting templates (as shown in Figures 12 and 13) for these models. Instead, to elicit the best possible performance from each model and establish the strongest baseline, we followed this procedure:
  1. 1.
    We cloned the official public repository associated with each chart-specific model.
  2. 2.
    We utilized their provided, out-of-the-box inference scripts and default model weights without modification.
  3. 3.
    We fed the images and questions from our CRBench test set into these scripts.
This methodology ensures that the results reported in the main body of the manuscript represent each model performing under its own optimal, author-intended conditions. This prevents any potential performance degradation that might arise from imposing a foreign prompt structure and guarantees that we are comparing ChartReasoner against the most capable version of each specialized model.
Evaluation of ChartReasoner on Public Benchmarks
To validate the generalization of ChartReasoner’s enhanced reasoning capabilities, we evaluated it against several standard public benchmarks. The following subsections detail the exact procedure used for each.
CharXiv
For the CharXiv benchmark, we utilized the official code and evaluation scripts available in their repository. Our evaluation process involved adapting their provided inference pipeline. We integrated our locally-deployed ChartReasoner (served via vLLM) as the model backend, replacing the default model call. All other components, including data loading, pre-processing, and the final scoring scripts for both the reasoning (R) and descriptive (D) tasks, were kept identical to the original setup.
ChartBench
Our evaluation on ChartBench followed a similar protocol. We used the complete evaluation pipeline provided in the official ChartBench repository. Predictions for each test instance were generated by passing the questions to our ChartReasoner vLLM endpoint. The resulting outputs were then fed directly into the official scoring script, which computes the overall accuracy based on their defined metrics and question categories.
ChartQAPro
The official repository for ChartQAPro provides a standalone evaluation script but does not include a full inference pipeline. Therefore, to benchmark ChartReasoner, we implemented a two-step process that carefully reconstructs the intended evaluation conditions: 1. Prediction Generation: We developed a custom script to generate predictions for the entire ChartQAPro test set. For each question, this script formatted a query for ChartReasoner that precisely replicated the direct prompt template described in the official ChartQAPro paper. 2. Scoring: The complete set of generated predictions was saved to a file. This file was then used as input for the official evaluation script provided by the ChartQAPro authors. The script automatically compares our model’s predictions against the ground-truth annotations to compute the final accuracy score.
Figure 6: An overview of sample question-answer pairs for various synthetic chart types within the CRBench dataset.
Figure 7: An overview of sample question-answer pairs for various real chart types within the CRBench dataset.
Figure 8: The Prompt in Distilling the CoT Data
Figure 9: Example 1 of CoT Dataset
Figure 10: Example 2 of CoT Dataset
Figure 11: Direct Prompting in CRBench Evaluation
Figure 12: Prompting for Optional Output of Reasoning Steps in CRBench Evaluation
Figure 13: Prompting for Forced Output of Reasoning Steps in CRBench Evaluation
Figure 14: Example 1 from CRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartReasoner-7B.
Figure 15: Example 2 from CRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartReasoner-7B.

### v2

@@
 Although Multimodal Large Language Models (MLLMs) have demonstrated increasingly impressive performance in chart understanding, most of them exhibit alarming hallucinations and significant performance degradation when handling non-annotated charts [^1] The non-annotated charts are those that require viewers to estimate values using the vertical/horizontal axis scale..
-Therefore, a question arises: Do MLLMs really understand the charts?
-Since a human is capable of understanding charts and estimating the values by visual reasoning, we first carefully establish a comprehensive Chart Reasoning Benchmark (CRBench) to rigorously evaluate the visual reasoning abilities of MLLMs on non-annotated charts.
-We argue that MLLMs are primarily relying on recognition rather than reasoning to interpret the charts.
-To steer MLLMs to reasonable chart understanding, we propose ChartReasoner that mimics human behavior by grounding their estimation in chart understanding.
-Extensive results on the proposed CRBench show that ChartReasnoner-3B/7B achieves superior performance in chart reasoning, even compared to GPT-4o and Gemini-2.5-Flash.
-More importantly, ChartReasnoner also demonstrates the visual reasoning abilities in general chart comprehension on public benchmarks, leading to significant performance gains and enabling MLLMs to rationally understand the charts.
+We argue that current MLLMs rely largely on visual recognition rather than visual reasoning to interpret the charts, and visual estimation of numerical values is one of the most fundamental capabilities in chart understanding that require complex visual reasoning.
+To prove this, we introduce ChartVRBench, a benchmark meticulously designed to isolate and evaluate visual reasoning ability in chart understanding.
+Furthermore, we propose ChartVR-3B/7B trained with a novel Visual Reasoning Reinforcement Finetuning (VR-RFT) strategy to strengthen genuine chart visual reasoning abilities.
+Extensive experiments show that ChartVR achieves superior performance on ChartVRBench, outperforming even powerful proprietary models.
+Moreover, the visual reasoning skills cultivated by the proposed VR-RFT demonstrate strong generalization, leading to significant performance gains across a diverse suite of public chart understanding benchmarks.
 The code and dataset will be publicly available upon publication.
-Introduction
-Multimodal Large Language Models (MLLMs) (Bai et al.
-2025; Comanici et al.
-2025; OpenAI et al.
-2024a; Lu et al.
-2024) now play a pivotal role in the field of Artificial Intelligence, particularly for understanding complex visual data.
-These models have demonstrated a remarkable ability to process charts, analyze their content, provide insightful explanations, and achieve competitive performance on many existing chart benchmarks (Wang et al.
-2024; Masry et al.
-2022; Xu et al.
-2024b; Masry et al.
-2025a; Xia et al.
-2025).
-However, we discover an intriguing phenomenon: When specific numerical annotations are missing in the charts, in other words, when handling non-annotated charts, the MLLMs’ propensity for hallucination increases dramatically—-a failure exemplified in Figure 1, where a model defaults to hallucination in non-annotated charts while our proposed method succeeds through multi-step reasoning.
+1 Introduction
+Multimodal Large Language Models (MLLMs) (bai2025qwen2; comanici2025gemini; hurst2024gpt; lu2024ovis) now play a pivotal role in the field of Artificial Intelligence, particularly for understanding complex visual data.
+These models have demonstrated a remarkable ability to process charts, analyze their content, provide insightful explanations, and achieve competitive performance against existing chart benchmarks (wang2024charxiv; masry2022chartqa; xu2023chartbench; masry2025chartqapro; xia2024chartx).
+Estimating numerical values from charts is a fundamental capability in chart understanding that involves interpreting visual representations to extract or approximate the underlying numbers.
+The core principle is to understand the mapping between the visual elements (e.g., the position, length, or angle of a mark) on the chart and the data scale it represents.
+However, when specific numerical annotations are missing from the chart, the propensity of MLLMs to hallucination increases dramatically (xu2023chartbench), as exemplified in Figure 1.
 This leads us to a fundamental question: Do MLLMs really understand the charts?
+This failure suggests that current MLLMs excel at recognizing about textual content within charts but struggle profoundly with reasoning from their underlying visual geometry.
+We argue that it stems from the fundamental reliance of MLLMs on textual recognition over genuine visual reasoning.
+To systematically diagnose this core ability, we introduce the Chart Visual Reasoning Benchmark (ChartVRBench), which is meticulously designed to isolate numerical value estimation on non-annotated charts, forcing models to move beyond textual recognition.
+The evaluation reveals that not only open-source MLLMs (bai2025qwen2; zhu2025internvl3; lu2024ovis) face performance degradation, but even powerful close-source MLLMs, such as GPT-4o (hurst2024gpt) and Gemini-2.5-Flash (comanici2025gemini), also struggle significantly with ChartVRBench.
 Figure 1: Illustration of the visual reasoning deficit in MLLMs when processing non-annotated charts.
-A “standard Initial Model”, limited by its underdeveloped visual reasoning capacity, often resorts to guessing and fails.
-In contrast, our “Optimized Model” (ChartReasoner) executes a deliberate, human-like reasoning chain—identifying the target, locating data based on visual scales, and forming a conclusion—to successfully estimate the value.
-To investigate this issue, we elaborately introduce a Chart Reasoning Benchmark (CRBench), a novel benchmark that specifically requires visual reasoning for chart understanding, focused on the task of numerical estimation on non-annotated charts.
-Our goal was to rigorously evaluate the visual reasoning capabilities of current MLLMs on chart understanding and explore whether they truly comprehend the visual clues in the charts.
-The evaluation reveals that not only open-source MLLMs (Bai et al.
-2025; Zhu et al.
-2025; Lu et al.
-2024) face performance degradation, but even powerful close-source MLLMs, such as GPT-4o (OpenAI et al.
-2024a) and Gemini-2.5-Flash (Comanici et al.
-2025), also struggle significantly with CRBench.
-Consequently, we argue that most MLLMs are much more relying on visual recognition capabilities, e.g., Optical Character Recognition (OCR), to read the charts, rather than employing genuine visual reasoning to truly understand them.
-Inspired by the success of Reinforcement Learning (RL) in enhancing textual reasoning in domains like mathematics and coding (DeepSeek-AI et al.
-2025; OpenAI et al.
-2024b), we introduce ChartReasoner, a new model forged through a novel Reinforcement Finetuning (RFT) strategy.
-This two-stage framework is designed not just to recognize the charts, but also understand the charts via visual reasoning.
-It begins with a “cold start” through Chain-of-Thought Supervised Fine-Tuning (CoT-SFT) to activate a structured thinking paradigm within the MLLM.
-This initial ability is then dramatically amplified in the second stage, where the Group Relative Policy Optimization (GPRO) (Shao et al.
-2024) training method selectively reinforces correct reasoning pathways and penalizes flawed logic.
-The experimental results demonstrate that this strategy significantly enhances the visual reasoning capabilities of ChartReasoner, achieving superior performance on the CRbench.
-Besides, we discovered that the foundational visual reasoning skills cultivated by our method generalize broadly, leading to substantial performance gains across a diverse suite of public chart understanding benchmarks.
-CharReasoner exhibits highly potential in moving towards a more genuine form of chart comprehension.
-The main contributions are summarized as follows:
-  • We introduce CRBench, a distinctive benchmark designed to isolate and evaluate genuine visual reasoning capability in chart understanding.
+A standard model, limited by its underdeveloped visual reasoning capacity, often resorts to guessing and fails.
+In contrast, our ChartVR executes a deliberate, human-like reasoning chain—identifying the target, locating data based on visual scales, and forming a conclusion—to successfully estimate the value.
+Moreover, inspired by the success of Reinforcement Learning (RL) in enhancing textual reasoning for mathematics and coding (guo2025deepseek; jaech2024openai; tan2025reason; huang2025vision), we propose ChartVR, a series of MLLMs forged with a novel Visual Reasoning Reinforcement Finetuning (VR-RFT) strategy to strengthen genuine chart visual reasoning abilities.
+The first stage, Visual Reasoning Activation, uses a Chain-of-Thought Supervised Finetuning (CoT-SFT) (liu2023visual) to compel the model to externalize a step-by-step analysis of the chart’s visual components.
+This forces the model to learn an explicit protocol for geometric interpretation, such as locating axes and grounding queries to graphical marks, thereby forming the structural foundation of its visual reasoning capability.
+Building on this, the second stage, Visual Reasoning Generalization, employs Group Relative Policy Optimization (GRPO) (shao2024deepseekmath) to further refine this process.
+By training on a curated dataset of ambiguous samples where the initial model’s judgment is inconsistent, we force it to make finer perceptual discriminations.
+This training process is guided by a novel continuous accuracy reward function with a quadratic formulation, providing a dense signal directly proportional to the accuracy of the visual estimation.
+In summary, these stages steer ChartVR to a robust, generalizable visual reasoning capability for charts.
+The extensive experiments demonstrate that ChartVR achieves superior performance on ChartVRBench, even comparable to powerful proprietary models like Gemini-2.5-Flash (comanici2025gemini).
+More importantly, we demonstrate that the foundational skill cultivated by our method is highly generalizable.
+ChartVR exhibits significant performance gains across a diverse suite of public, multi-task chart understanding benchmarks (wang2024charxiv; xu2023chartbench; masry2025chartqapro), proving the effectiveness of our approach in building more rational and reliable MLLMs for chart comprehension.
+The main contributions of this work are summarized as follows:
+  • We introduce ChartVRBench, a distinctive benchmark designed to isolate and evaluate genuine visual reasoning capability in chart understanding.
 It overcomes the limitations of prior work by focusing exclusively on numerical estimation tasks, thus disentangling reasoning from text recognition.
-  • We propose ChartReasoner, a powerful MLLM with significantly enhanced visual reasoning capabilities for chart understanding.
-It achieves excellent performance on our challenging CRBench benchmark, compared with chart-specific and general MLLMs, even surpassing powerful proprietary models like GPT-4o and Gemini-2.5-Flash.
+  • We propose ChartVR, a series of MLLMs with significantly enhanced visual reasoning capabilities for chart understanding.
+It achieves excellent performance on our challenging ChartVRBench, compared with chart-specific and general MLLMs, even surpassing powerful proprietary models like Gemini-2.5-Flash.
   • We demonstrate that the visual reasoning ability cultivated by our method is foundational and highly generalizable.
-ChartReasoner is not confined to the specific numerical estimation task, but achieves substantial performance gains across a diverse suite of public, multi-task chart understanding benchmarks.
-Related Work
-Chart Understanding
-The advent of MLLMs, encompassing both powerful open-source and close-source (Comanici et al.
-2025; OpenAI et al.
-2024a; Team et al.
-2025), has laid a foundational groundwork for automated chart understanding.
-Driven by large-scale visual instruction tuning (Liu et al.
-2023), this progress has spurred the development of numerous chart-specific fine-tuned models (Han et al.
-2023; Zhang et al.
-2024; Masry et al.
-2024, 2025b) tailored to enhance performance on these specialized tasks.
-However, the predominant training paradigm for these models relies on SFT (Liu et al.
-2023; Wei et al.
-2022), which, as we argue, can lead to superficial mimicry rather than genuine reasoning.
-In contrast, our work explores a paradigm based on RL (Ouyang et al.
-2022) to cultivate a deeper, more robust form of chart understanding.
-Figure 2: The training paradigm of ChartReasoner and the data distribution of CRBench.
-A: ChartReasoner leverages a two-stage RFT strategy.
+ChartVR is not confined to the specific numerical estimation task, but achieves substantial performance gains across a diverse suite of public, multi-task chart understanding benchmarks.
+2 Related Work
+2.1 Chart Understanding Benchmarks
+A suite of benchmarks has been developed to evaluate the chart comprehension capabilities of MLLMs.
+Early benchmarks, such as ChartQA (masry2022chartqa) and PlotQA (methani2020plotqa), primarily focused on descriptive tasks.
+More recently, benchmarks like CharXiv (wang2024charxiv), ChartQAPro (masry2025chartqapro), and ChartMuseum (tang2025chartmuseum) have raised the bar by incorporating complex questions and diverse, real-world charts.
+While these works encompass a wide range of tasks, they often conflate general reasoning with the core challenge of visual interpretation.
+The most related work to ours is ChartBench (xu2023chartbench); while it also focuses on non-annotated charts, it is composed of mostly synthetic data with limited visual diversity.
+Similarly, recent work by mukhopadhyay2024unraveling revealed critical flaws in the consistency and robustness of MLLMs but stopped short of attributing these shortcomings to a fundamental deficit in visual reasoning.
+We argue this deficit—the core skill of visual reasoning in a chart’s geometry, such as numerical value estimation—remains largely untested.
+Our ChartVRBench is specifically designed to isolate and evaluate this crucial visual reasoning capability.
+2.2 Chart Understanding with MLLMs
+Many general-purpose MLLMs, such as gpt-4o (hurst2024gpt), Gemini-2.5 Series (comanici2025gemini), and Qwen (bai2025qwen2), are increasingly applied to chart understanding tasks.
+In parallel, the development of specialized Chart MLLMs has been rapid, with many models like ChartLlama (han2023chartllama) and ChartGemma (masry2024chartgemma).
+However, their development has largely depended on SFT, a paradigm that, as we argue, tends to cultivate superficial recognition at the expense of genuine reasoning.
+Recognizing this, a recent wave of models (chen2025chartr1; masry2025bigcharts), have incorporated RL to enhance complex, multi-step reasoning.
+While these RL-based approaches represent a significant step forward, their training objectives often prioritize the final accuracy of text-heavy queries, which can leave the foundational skill of visual grounding underdeveloped.
+In contrast, our ChartVR is specifically designed to address this fundamental layer.
+Its RFT framework is meticulously crafted to cultivate the core ability to reason directly from visual geometry, aiming to develop a genuine visual reasoning capability rather than optimizing the textual reasoning that typically follows.
+2.3 Reasoning in Chart Understanding
+Reinforcement Learning (RL) has been successfully employed to enhance the reasoning abilities of Large Language Models (LLMs), allowing them to move beyond the static data distributions of SFT (ouyang2022training).
+By learning from reward feedback, models have shown significant improvements in complex domains like mathematics and coding (guo2025deepseek; shao2024deepseekmath).
+Inspired by this success, several works have begun to apply similar RL-based paradigms to MLLMs (feng2025video; tan2025reason; huang2025vision), activating their visual reasoning on tasks like visual counting and spatial transformation.
+Building on these advancements, our work adapts this powerful paradigm to the specialized domain of chart understanding.
+Figure 2: The training paradigm of ChartVR and the data distribution of ChartVRBench.
+A: ChartVR leverages a two-stage RFT strategy.
 Stage 1 activates the model’s reasoning abilities via SFT on CoT data, while Stage 2 uses GRPO with a multi-component reward system to reinforce correct chart understanding.
-B: The composition of CRBench, detailing the distribution of seven chart types across both synthetic and real data sources.
-Methods Synthetic Charts Real Charts Overall
-Box Area Radar Scatter Bar Line Combo Bar Line Combo
-Human 74.72 56.70 83.97 77.90 90.10 93.71 77.27 80.36 65.22 52.00 78.64
-Open-source Models
-InternVL3-2B (Zhu et al. 2025) 31.87 7.32 5.77 23.19 51.40 45.70 32.98 39.29 33.04 20.80 31.59
-Qwen2.5-VL-3B (Bai et al. 2025) 49.45 15.24 19.23 52.74 74.58 82.45 62.48 67.86 54.78 48.80 57.68
-Ovis1.6-llama3.2-3B (Lu et al. 2024) 16.48 2.44 10.90 14.22 17.04 12.25 11.52 12.50 13.04 9.60 12.47
-Gemma-3-4B (Team et al. 2025) 15.38 3.05 10.90 10.28 8.38 8.94 6.28 5.36 9.57 8.80 8.32
-Qwen2.5-VL-7B (Bai et al. 2025) 72.53 28.05 26.92 66.08 84.08 86.42 71.03 63.39 60.87 47.20 66.25
-InternVL3-8B (Zhu et al. 2025) 53.85 19.51 14.10 43.33 62.29 61.92 48.52 58.93 55.65 42.40 42.40
-Close-source Models
-GPT-4o (OpenAI et al. 2024a) 32.97 15.24 12.82 25.82 22.35 25.17 19.90 15.18 26.96 16.80 21.69
-Gemini-2.5-Flash (Comanici et al. 2025) 72.53 32.93 11.54 59.96 70.11 71.19 55.50 53.57 46.96 37.60 55.32
-Chart-Specific Models
-ChartGemma-3B (Masry et al. 2025b) 10.99 10.98 7.05 21.44 42.74 32.78 24.43 37.50 42.61 28.80 26.74
-TinyChart-3B (Zhang et al. 2024) 13.19 7.93 7.69 25.16 56.15 54.30 36.65 57.14 40.87 34.40 35.83
-ChartInstruct-7B (Masry et al. 2024)
-10.99 1.22 4.49 16.63 35.47 16.56 18.85 51.79 45.22 18.40 20.91
-ChartVLM-7.3B (Xia et al. 2025) 9.89 10.37 7.69 10.28 70.39 45.36 32.81 50.00 54.78 35.20 33.63
-ChartLlama-13B (Han et al. 2023) 10.99 1.83 5.77 5.25 3.35 4.30 4.01 8.04 9.57 7.20 5.01
-Table 1: Performance evaluation on CRBench with direct answering, covering a range of open-source, close-source, and chart-specific models.
-This table shows the results when models are prompted to directly output the final answer without intermediate reasoning steps.
-A suite of chart benchmarks, including ChartQA (Masry et al.
-2022), PlotQA (Methani et al.
-2020), and others (Xu et al.
-2024b; Han et al.
-2023; Wang et al.
-2024), have been proposed to evaluate MLLMs on diverse chart tasks.
-While instrumental, their task design often prioritizes language-level reasoning (Lewkowycz et al.
-2022) (e.g., calculating averages or summarizing) over fundamental visual reasoning (Johnson et al.
-2017; Hudson and Manning 2019).
-The core skill of grounding reasoning in a chart’s geometry, such as estimating values from axes on non-annotated charts, remains largely untested.
-Our CRBench is specifically designed to isolate and evaluate this crucial visual reasoning capability.
-Visual Reasoning
-RL has been successfully employed to significantly enhance the reasoning abilities of Large Language Models (LLMs) (DeepSeek-AI et al.
-2025), particularly in complex domains like mathematics and coding (Shao et al.
-2024).
-Inspired by this success in language-only models, several works have begun to apply similar RL-based paradigms to MLLMs (Feng et al.
-2025; Tan et al.
-2025; Huang et al.
-2025), successfully activating their visual reasoning capabilities on various challenging tasks (e.g., visual counting, spatial transformation).
-Building on these advancements, our work adapts and applies an RL-based framework to the specialized domain of chart understanding.
-Chart Reasoning Benchmark
-We introduce Chart Reasoning Benchmark (CRBench), a comprehensive, multi-domain, and reasoning-centric benchmark designed to rigorously assess the visual interpretation capabilities of MLLMs on charts that lack explicit numerical annotations.
+B: The composition of ChartVRBench, detailing the distribution of seven chart types across both synthetic and real data sources.
+3 ChartVRBench
+We introduce Chart Visual Reasoning Benchmark (ChartVRBench), a comprehensive, multi-domain, and reasoning-centric benchmark designed to rigorously assess the visual interpretation capabilities of MLLMs on charts that lack explicit numerical annotations.
 Engineered to move beyond simple OCR-dependent tasks, the benchmark comprises a total of 2,453 question-answer pairs.
@@
 The benchmark provides extensive coverage across seven primary chart types, including bar, line, scatter, and combo charts, with a detailed breakdown of the data distribution shown in Figure 2.
-This structural diversity is complemented by thematic breadth, with data spanning 38 distinct topics, including finance, healthcare, and technology, etc.
+This structural diversity is complemented by thematic breadth, with data spanning 38 distinct topics, including finance, healthcare, and technology.
 This dual emphasis on structural and thematic variety ensures a rigorous evaluation, mitigating the risk of models overfitting to specific chart formats or familiar domains.
-Data Curation
-Our data generation process integrates automated synthesis with manual curation to ensure both scale and quality.
+While existing benchmarks have significantly advanced the field, they predominantly focus on general high-level Question Answering (QA), where visual reasoning is often conflated with textual extraction (OCR) and logical reasoning.
+ChartVRBench fills a critical gap by strictly isolating the visual reasoning capability of numerical value estimation on non-annotated charts, preventing models from relying on text recognition shortcuts.To clearly demonstrate how our benchmark compares to contemporary works, we present a feature-wise comparison in Table 1.
+Table 1: Comparison between ChartVRBench and existing representative chart QA benchmarks.
+Symbols: ✓ Fully Supported / High Quality; $\triangle$ Partially Supported / Mixed; $\times$ Not Supported / Low Quality.
+Feature
+EvoChart
+huang2025evochart
+ChartBench
+xu2023chartbench
+CharXiv
+wang2024charxiv
+ChartQAPro
+masry2025chartqapro
+ChartMuseum
+tang2025chartmuseum
+Ours
+Real-World Charts $\times$
+Broad Topic Coverage $\times$ $\times$ $\triangle$ $\triangle$ $\triangle$
+Non-Annotated Charts $\times$ $\times$ $\times$ $\times$
+Isolates Visual Reasoning $\times$ $\triangle$ $\times$ $\times$
+3.1 Data Curation
 Synthetic Chart Generation.
-Our synthetic chart generation process is partially adapted from the Code-as-Intermediary Translation (CIT) methodology proposed by (He et al.
-2024), where executable plotting code serves as the ground truth for each chart.
-The process begins with a curated set of seed scripts, which are then programmatically diversified using Self-Instruct (Wang et al.
-2023) and Evol-Instruct (Xu et al.
-2024a) techniques to generate a vast library of visually complex charts.
+Our synthetic chart generation process is partially adapted from the Code-as-Intermediary Translation (CIT) methodology proposed by he2024distill, where executable plotting code serves as the ground truth for each chart.
+The process begins with a curated set of seed scripts, which are then programmatically diversified using Self-Instruct (wang2022self) and Evol-Instruct (xu2024wizardlm) techniques to generate a vast library of visually complex charts.
 A critical constraint is the deliberate omission of numerical labels on data points, ensuring that every chart necessitates visual estimation.
 To maximize yield, a self-repair mechanism leverages an LLM to debug and correct any code that fails during execution.
-Following an automated visual fidelity check by a multimodal language model, the entire collection of synthesized data underwent a final, rigorous human review.
+Following an automated visual fidelity check by a MLLM, the entire collection of synthesized data underwent a final, rigorous human review.
 This manual verification step served to confirm the high quality of the chart images and the accuracy of their corresponding question-answer pairs.
@@
 Every MLLM-generated pair then underwent a final round of human verification and refinement to guarantee the accuracy and relevance of both the question and its ground-truth answer.
-Evaluation Protocol
+3.2 Evaluation Protocol
 Standard exact-match accuracy is ill-suited for value estimation from non-annotated charts, as it fails to account for the slight perceptual ambiguity inherent in the task, even for human observers.
@@
 This protocol ensures that our evaluation is both rigorous and fairly aligned with human-level interpretive capabilities, rewarding models for precise visual reasoning rather than penalizing them for minor, human-like estimation variance.
-Methods Synthetic Charts Real Charts Overall
-Box Area Radar Scatter Bar Line Combo Bar Line Combo
-Open-source Models
-InternVL3-2B (Zhu et al. 2025) 25.27 8.54 9.62 25.16 51.68 43.05 34.55 43.75 34.78 32.00 32.98
-Qwen2.5-vl-3B (Bai et al. 2025) 46.15 14.63 17.95 52.30 72.35 80.79 62.65 66.96 53.04 46.40 56.62
-Ovis1.6-llama3.2-3B (Lu et al. 2024) 12.09 3.66 7.69 14.66 13.13 10.93 11.69 8.04 16.52 12.00 11.66
-Gemma-3-4B (Team et al. 2025) 9.89 3.05 12.82 11.82 9.22 12.25 6.81 8.04 12.17 7.20 9.34
-Qwen2.5-vl-7B (Bai et al. 2025) 69.23 20.12 19.87 61.93 74.02 83.77 67.71 49.11 53.04 46.40 60.74
-InternVL3-8B (Zhu et al. 2025) 36.73 12.80 12.18 39.17 43.58 47.68 36.82 38.39 46.09 34.40 36.73
-Close-source Models
-GPT-4o (OpenAI et al. 2024a) 28.57 12.20 11.54 25.61 21.23 27.15 18.15 13.39 26.96 18.40 20.87
-Gemini-2.5-Flash (Comanici et al. 2025) 68.13 25.61 7.69 61.93 72.07 75.17 55.85 49.11 52.17 39.20 55.77
-Chart-Specific Models
-ChartReasoner-3B (Ours) 58.24 21.95 22.44 56.24 79.61 86.42 66.14 65.18 54.78 48.80 61.27
-ChartReasoner-7B (Ours) 72.53 31.71 32.05 65.43 82.96 84.44 71.55 65.18 66.09 53.60 67.06
-Table 2: Comparison of ChartReasoner with representative MLLMs on the porposed CRBench, using Chain-of-Thought (CoT) prompting to elicit the models’ intermediate reasoning steps.
-The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
-ChartReasoner
-We propose ChartReasoner, a MLLM designed to perform visual reasoning for better visual understanding on non-annotated charts.
+4 ChartVR
+We propose ChartVR, a series of MLLMs designed to perform visual reasoning for better visual understanding on non-annotated charts.
 We formally define this task as follows: given a chart image $I$, and a corresponding textual question $Q$, the goal is to derive a numerical answer $A$ with a reasoning procedure $R$.
@@
 The fundamental challenge lies in interpreting non-annotated charts, which requires the model to reason about geometric structures (e.g., axes, scales, positions) to infer values, rather than simply extracting them via text recognition.
-While the mapping $\mathcal{F}$ inherently performs this reasoning, producing an explicit reasoning chain $R$ that explains the derivation is an optional, secondary objective.
-To address this challenge, we propose a two-stage Reinforcement Finetuning (RFT) strategy.
-This approach is designed to first instill a robust, human-like reasoning framework and then meticulously refine the model’s numerical precision.
-As illustrated in Figure 2, the RFT pipeline consists of two sequential stages: (1) Reasoning Activation via SFT, which uses supervised fine-tuning to teach the model the structure of reasoning, followed by (2) Reasoning Generalization via RL, which uses reinforcement learning to improve the accuracy and generalizability.
-Stage 1: Reasoning Activation via SFT
+To address this challenge, we propose a novel two-stage Reinforcement Finetuning (RFT) framework.
+This approach is designed to first instill a robust, human-like reasoning framework and then meticulously refine the model’s numerical e precision.
+As illustrated in Figure 2, the RFT pipeline consists of two sequential stages: (1) Visual Reasoning Activation, which uses supervised fine-tuning to teach the model the structure of reasoning, followed by (2) Visual Reasoning Generalization, which uses reinforcement learning to improve the accuracy and generalizability.
+4.1 Stage 1: Visual Reasoning Activation
 The initial stage of our pipeline aims to establish a foundational reasoning paradigm.
 Instead of having the model directly guess an answer, we teach it to adopt a structured, step-by-step thought process that mirrors human analysis.
-To achieve this, we first generate a high-quality dataset by distilling detailed Chain-of-Thought (CoT) processes from a MLLM with emerging reasoning abilities.
-Fine-tuning our base MLLM on these 68,500 synthesized samples teaches the model to emulate a human’s analytical approach, guiding it to identify and utilize critical chart components—such as axes, scales, and legends—before arriving at a final answer.
+To achieve this, we fine-tune our base model on a high-quality dataset of 43k samples generated by distilling detailed Chain-of-Thought (CoT) processes from an advanced MLLM (see Appendix B.1 for details).
+This CoT-SFT process systematically teaches the model to move beyond direct answer prediction and instead adopt a structured analytical approach: first identifying and utilizing critical chart components—such as axes, scales, and legends—and then using them to derive a final answer.
 Formally, we employ SFT on this dataset.
 Each data instance is a tuple $(x,q,r,a)$, where $x$ is the chart image, $q$ is the question, $r$ is the intermediate reasoning chain, and $a$ is the final answer.
-The training objective is to maximize the log-likelihood of the model generating the complete sequence $y$ (the concatenation of $r$ and $a$) given the image $x$ and question $q$:
+The training objective is to minimize the negative log-likelihood of the model generating the complete sequence $y$ (the concatenation of $r$ and $a$) given the image $x$ and question $q$:
 $\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,q,r,a)\sim\mathcal{D}}\sum_{t=1}^{|y|}\log\pi_{\theta}(y_{t}|x,q,y_{<t})$ (1)
@@
 The resulting fine-tuned model, denoted as $\pi_{\text{SFT}}$, learns a robust template for visual reasoning and serves as the starting point for the next stage.
-Stage 2: Reasoning Generalization via RL
-Building on the reasoning foundation from Stage 1, the second stage focuses on enhancing the model’s precision and reliability for the numerical estimation task.
-For this, we use a smaller, high-signal dataset of 3,400 samples curated to target the model’s specific weaknesses.
-These samples are identified by selecting problems where the SFT-tuned model exhibits ”stochastic correctness”—that is, problems it can solve but not consistently.
+4.2 Stage 2: Visual Reasoning Generalization
+Building on the visual reasoning foundation from Stage 1, the second stage focuses on enhancing the model’s precision and reliability for the numerical estimation task.
+For this, we use a smaller, high-signal dataset of 3.4k samples curated to target the model’s specific weaknesses.
+These samples are identified by selecting problems where the SFT-tuned model exhibits “stochastic correctness”—that is, problems it can solve but not consistently (see Appendix B.3 for details).
 By training on these borderline cases with higher-resolution images, we force the model to refine its visual interpretation skills.
-We employ GRPO (Shao et al.
-2024), an efficient and scalable reinforcement learning algorithm, to fine-tune the policy $\pi_{\text{SFT}}$.
-Unlike traditional algorithms like PPO (Schulman et al.
-2017), GRPO forgoes a computationally expensive value network and instead calculates relative advantages by comparing rewards within a group of sampled responses.
+We employ GRPO (shao2024deepseekmath), an efficient and scalable reinforcement learning algorithm, to fine-tune the policy model $\pi_{\text{SFT}}$.
+Unlike traditional algorithms like PPO (schulman2017proximalpolicyoptimizationalgorithms), GRPO forgoes a computationally expensive value network and instead calculates relative advantages by comparing rewards within a group of sampled responses.
 For each input $(x,q)$, we sample a group of $G$ candidate answers $\{a_{1},a_{2},\dots,a_{G}\}$ from the current policy $\pi_{\beta}$.
@@
 The policy is then updated to increase the probability of actions with positive advantages, while a KL-divergence penalty against the reference model $\pi_{\text{SFT}}$ ensures stable training.
-Reward Function Design
+4.3 Reward Function Design
 The effectiveness of our RL stage hinges on a well-designed reward function. Our function $R(a_{i})$ is a composite of two components, targeting both response structure and numerical accuracy:
@@
 To encourage interpretable and well-structured outputs, we provide a binary format reward, $R_{\text{format}}$.
-The model receives a reward of 1 if its response strictly adheres to our predefined template—reasoning enclosed in <think></think> tags and the final answer in <answer></answer> tags—and 0 otherwise.
-Accuracy Reward.
+The model receives a reward of 1 if its response strictly adheres to our predefined template, where reasoning is enclosed in <think></think> tags and final answer in <answer></answer> tags, and 0 otherwise.
+Continuous Accuracy Reward.
 To overcome the sparse signal from a simple correct/incorrect binary reward, we introduce a continuous accuracy reward, $R_{\text{acc}}$.
@@
 $d_{\text{rel}}=\frac{\lvert A_{\text{pred}}-A_{\text{gt}}\rvert}{\lvert A_{\text{gt}}\rvert}$ (4)
-Based on a human-calibrated tolerance threshold of $\varepsilon=0.02$, we define the reward using a piecewise quadratic function that smoothly decays from 1 to 0:
-$R_{\text{acc}}(d_{\text{rel}})=\begin{cases}\left(1-\frac{d_{\text{rel}}}{\varepsilon}\right)^{2}&\text{if }d_{\text{rel}}<\varepsilon\\ 0&\text{if }d_{\text{rel}}\geq\varepsilon\end{cases}$ (5)
+Then, we define the reward using a piecewise quadratic function that smoothly decays from 1 to 0:
+$R_{\text{acc}}(d_{\text{rel}})=\begin{cases}\left(1-\frac{d_{\text{rel}}}{\tau}\right)^{2}&\text{if }d_{\text{rel}}<\tau\\ 0&\text{if }d_{\text{rel}}\geq\tau\end{cases}$ (5)
+We empirically set $\tau=0.02$ based on a human-calibrated tolerance threshold.
+For the specific case where the ground truth $A_{\text{gt}}$ is zero, because it is difficult to quantize the relative deviations, the accuracy reward falls back to exact match, assigning a value of 1 when $A_{\text{gt}}=A_{\text{pred}}$ and 0 otherwise.
 We employ the quadratic formulation for two critical reasons.
@@
 This behavior encourages the model to make fine-grained improvements when it is already close to the correct answer, while strongly penalizing larger, unacceptable deviations.
-Figure 3: Qualitative analysis on CRBench.
+Table 2: Comparison of ChartVR with representative MLLMs on the proposed ChartVRBench. The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
+Methods Synthetic Charts Real Charts Overall
+Box Area Radar Scatter Bar Line Combo Bar Line Combo
+Human Evaluation 94.51 43.29 88.46 91.24 96.65 97.68 90.92 84.82 84.35 65.60 87.57
+Open-source Models
+InternVL3-2B (zhu2025internvl3) 25.27 8.54 9.62 25.16 51.68 43.05 34.55 43.75 34.78 32.00 32.98
+Qwen2.5-vl-3B (bai2025qwen2) 46.15 14.02 17.95 51.42 72.91 81.13 62.83 66.96 54.78 45.60 56.62
+Ovis1.6-llama3.2-3B (lu2024ovis) 12.09 3.66 7.69 14.66 13.13 10.93 11.69 8.04 16.52 12.00 11.66
+Gemma-3-4B (team2025gemma) 9.89 3.05 12.82 11.82 9.22 12.25 6.81 8.04 12.17 7.20 9.34
+Qwen2.5-vl-7B (bai2025qwen2) 70.33 21.34 19.23 61.93 73.74 85.43 68.41 49.11 56.52 48.00 61.39
+InternVL3-8B (zhu2025internvl3) 36.73 12.80 12.18 39.17 43.58 47.68 36.82 38.39 46.09 34.40 36.73
+Close-source Models
+GPT-4o (hurst2024gpt) 28.57 12.20 11.54 25.61 21.23 27.15 18.15 13.39 26.96 18.40 20.87
+Gemini-2.5-Flash (comanici2025gemini) 68.13 25.61 7.69 61.93 72.07 75.17 55.85 49.11 52.17 39.20 55.77
+Chart-Specific Models
+ChartGemma-3B (masry2024chartgemma) 10.99 10.98 7.05 21.44 42.74 32.78 24.43 37.50 42.61 28.80 26.74
+TinyChart-3B (zhang2024tinychart) 13.19 7.93 7.69 25.16 56.15 54.30 36.65 57.14 40.87 34.40 35.83
+ChartInstruct-7B (masry2024chartinstruct) 10.99 1.22 4.49 16.63 35.47 16.56 18.85 51.79 45.22 18.40 20.91
+ChartVLM-7.3B (xia2024chartx) 9.89 10.37 7.69 10.28 70.39 45.36 32.81 50.00 54.78 35.20 33.63
+ChartLlama-13B (han2023chartllama) 10.99 1.83 5.77 5.25 3.35 4.30 4.01 8.04 9.57 7.20 5.01
+Bespoke-MiniChart-7B (liyan2025minichart) 72.53 26.83 25.00 66.74 86.87 89.40 75.04 69.64 62.61 45.60 68.16
+Chart-R1 (7B) (chen2025chartr1) 83.52 26.22 25.64 66.83 85.47 82.45 68.94 61.61 59.13 44.80 65.72
+ChartVR-3B (Ours) 63.74 24.39 19.87 57.33 82.68 87.75 69.46 58.93 57.39 45.60 62.74
+ChartVR-7B (Ours) 83.52 37.20 27.56 70.90 88.27 92.05 78.53 69.64 62.61 58.40 72.20
+5 Experiments
+5.1 Experimental Setups
+Implementation Details.
+The implementation was built upon the ModelScope SWIFT framework (zhao2025swift).
+We initialize our ChartVR models using the open-source Qwen2.5-VL series (bai2025qwen2) as a foundation.
+For inference, all models and benchmarks follow their provided settings where available, with results obtained from a single forward pass using a fixed random seed of 42 to ensure reproducibility.
+Additional details are available in Appendix C.
+Main Evaluation on ChartVRBench.
+Our primary evaluation is conducted on the proposed ChartVRBench to assess genuine visual reasoning capabilities and establish the superiority of our ChartVR model.
+On this benchmark, we compare our model against a comprehensive suite of baselines organized into three categories: open-source MLLMs, powerful close-source MLLMs and prominent chart-specific models.
+Generalization Study on Public Benchmarks.
+To evaluate the transferability of the skills learned via our RFT framework, we conduct a generalization study.
+Specifically, we benchmark our ChartVR model against the representative chart-specific models—ChartGemma (team2025gemma), TinyChart (zhang2024tinychart), ChartInstruct (masry2024chartinstruct), ChartVLM (xia2024chartx), and ChartLlama (han2023chartllama)—on a diverse set of public benchmarks.
+This suite includes the real-world datasets CharXiv (wang2024charxiv) and ChartQAPro (masry2025chartqapro), as well as the synthetic benchmark ChartBench (xu2023chartbench).
+This allows us to verify that our training methodology imparts a foundational reasoning ability that generalizes effectively to a variety of chart understanding tasks.
+5.2 Experimental Results
+Table 3: Performance of ChartVR compared to other chart-specific models on various public chart understanding benchmarks.
+All results are reproduced by the authors.
+Qwen2.5-VL baselines are listed below their corresponding ChartVR models with an arrow indicator.
+The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
+Models ChartVRBench CharXiv (R) ChartBench ChartQAPro
+ChartGemma-3B (masry2024chartgemma) 26.74 12.50 - 6.84
+TinyChart-3B (zhang2024tinychart) 35.83 8.30 - 13.25
+ChartInstruct-7B (masry2024chartinstruct) 20.91 8.80 - 4.88
+ChartVLM-7.3B (xia2024chartx) 33.63 - 12.06 -
+ChartLlama-13B (han2023chartllama) 5.01 14.20 21.30 -
+Bespoke-MiniChart-7B (liyan2025minichart) 68.16 41.40 44.19 34.74
+Chart-R1 (7B) (chen2025chartr1) 65.80 50.00 15.71 31.81
+ChartVR-3B (Ours) 62.74 33.40 26.35 28.03
+   $\hookrightarrow$ Qwen2.5-VL-3B (bai2025qwen2) 56.62 30.60 21.30 24.51
+ChartVR-7B (Ours) 72.20 43.40 45.34 41.79
+   $\hookrightarrow$ Qwen2.5-VL-7B (bai2025qwen2) 61.39 39.50 35.35 37.10
+Performance on ChartVRBench.
+The results, presented in Table 2, underscore the significant challenge that ChartVRBench poses to a wide range of MLLMs.
+The generally low scores across all categories—including powerful proprietary models like GPT-4o (20.87%) and Gemini-2.5-Flash (55.77%)—reveal a critical and widespread weakness in genuine visual reasoning.
+This difficulty stems from our benchmark’s design, which forces models to infer values from graphical geometry (e.g., axes and scales) rather than relying on OCR-based shortcuts common in other benchmarks.
+The particularly low score of GPT-4o has been double-checked and manually validated, which precisely indicates that many MLLMs lack the genuine visual reasoning ability, as we argue.
+Our proposed model, ChartVR, demonstrates a clear superiority in this demanding task.
+ChartVR-7B achieves an overall score of 72.20%, outperforming all other models, including the best open-source baseline, Qwen2.5-vl-7B (61.39%), and the strongest proprietary model, Gemini-2.5-Flash.
+Notably, even our smaller ChartVR-3B model (62.74%) surpasses most other models, highlighting the effectiveness of our training methodology.
+The performance is particularly strong on complex chart types like Real Combo charts, where ChartVR-7B (58.40%) dramatically outperforms the other models.
+Generalization on Public Benchmarks.
+As detailed in Table 3, ChartVR-7B exhibits exceptional generalization, achieving competitive results across the board.
+Specifically, on the reasoning-focused portion of CharXiv, our model achieves an improvement of 3.9% over the base model Qwen2.5-VL-7B.
+This success stems from our model’s core visual reasoning capability, which contrasts with other chart-specific models that rely on SFT and often fail to develop a generalizable reasoning capability.
+Table 4: Ablation study of training strategies for ChartVR-3B and ChartVR-7B models. The best and second-best scores are highlighted.
+Training Paradigm ChartVR-3B ChartVR-7B
+Synthetic Real Overall Synthetic Real Overall
+Zero-Shot 56.87 54.78 56.52 67.66 51.16 61.39
++CoT-SFT 36.55 34.86 40.77 55.45 45.74 54.06
++GRPO 56.59 56.82 56.62 66.44 54.83 64.78
++RFT 64.26 53.69 62.74 73.68 63.35 72.20
+Table 5: Ablation study of reward function components for Qwen2.5-VL-7B.
+Reward Component Score
+Format 61.84
+Cont. Acc 67.88
+Acc + Format 70.28
+Cont. Acc + Format 72.20
+5.3 Ablation Study
+Effectiveness of the RFT Training Paradigm.
+To systematically validate our training strategies, we conducted an comparative study comparing three paradigms: CoT-SFT, GRPO applied directly to the base model, and our full RFT framework.
+The results are summerized in Table 4.
+For ChartVR-7B, the base model scores 61.39%.
+Using CoT data as a ’Visual Reasoning Activation’ merely compels the model to adopt a visual reasoning pattern without imparting the underlying ability.
+Consequently, this mismatch leads to a performance degradation rather than an improvement.
+In contrast, the full RFT framework, which synergistically combines SFT with RL, achieves the most significant performance gain, reaching 72.20%.
+This demonstrates that our complete VR-RFT framework genuinely enhances the model’s visual reasoning ability, leading to its consistent and substantial outperformance over all other configurations.
+Impact of the Reward Function Design.
+We conducted an ablation study to isolate the contribution of each component in our reward function, with results presented in Table 5.
+The findings highlight a strong synergistic effect between enforcing a correct output structure and rewarding numerical precision.
+A model trained with only the Format reward achieves a score of 61.84, whereas combining it with our proposed Continuous Accuracy (Cont.
+Acc) reward boosts the score significantly to 72.20, indicating both components are crucial.
+Furthermore, the study validates the superiority of our continuous reward design over a standard binary alternative.
+A model using a simple binary accuracy reward (Acc + Format), which provides a sparse correct/incorrect signal, is clearly outperformed by our model using the continuous reward (Cont.
+Acc + Format).
+This demonstrates the effectiveness of our quadratic reward function, which provides a dense and informative learning gradient.
+By rewarding ”nearly correct” answers, it encourages the fine-grained improvements in visual estimation necessary for achieving higher precision.
+5.4 Case Study
+Figure 3: Qualitative analysis on ChartVRBench.
 GPT-4o and Qwen2.5-VL-7B exhibits hallucination in the answer and reasoning procedure, respectively.
-In contrast, our ChartReasoner-7B is able to produce a coherent and correct step-by-step reasoning process, leading to an accurate answer.
-Methods CRBench CharXiv (R) CharXiv (D) ChartBench ChartQAPro
-ChartGemma-3B (Masry et al. 2025b) 26.74 12.50 21.30 - 6.84
-TinyChart-3B (Zhang et al. 2024) 35.83 8.30 16.15 - 13.25
-ChartInstruct-7B (Masry et al. 2024) 20.91 8.80 21.40 - 4.88
-ChartVLM-7.3B (Xia et al. 2025) 33.63 - - 12.06 -
-ChartLlama-13B (Han et al. 2023) 5.01 14.20 19.23 21.30 -
-ChartReasoner-7B (Ours) 67.06 43.40 63.90 45.34 36.06
-Table 3: Performance of ChartReasoner compared to other chart-specific models on various public chart understanding benchmarks.
-The results of the compared methods are from the corresponding benchmarks.
-The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
-Experiments
-Experimental Setups
-Training Setting
-The implementation was built upon the ModelScope SWIFT framework (Zhao et al.
-2025) for all training procedures.
-We initialize ChartReasoner with Qwen2.5-VL (Bai et al.
-2025) and tune it with the dataset described in the previous section.
-All experiments were carried out on a server with 8 NVIDIA H800 GPUs.
-The methods evaluated on CRBench follow their default inference settings.
-All results were obtained from a single forward pass with a fixed random seed of 42.
-Please refer to the supplementary materials for more details.
-Training Paradigms and Baselines
-To systematically evaluate our training strategies, we conduct a comparative study in four distinct paradigms.
-  • CoT-SFT: Provides a reasoning “cold start” by fine-tuning the model on CoT dataset, with the reasoning data distilled from Qwen2.5-VL-32B (Bai et al. 2025).
-  • GRPO-Zero: An ablation study in which our GRPO algorithm is applied directly to the base model to assess the efficacy of reinforcement learning in isolation.
-  • RFT: Our proposed final paradigm starts from the CoT-SFT checkpoint and uses GRPO to further refine and generalize the model’s reasoning capabilities.
-To further validate the generalization and effectiveness of our training paradigms, we conduct experiments across a diverse set of base models.
-Specifically, we evaluate our methods on both the Qwen2.5-VL-Instruct (Bai et al.
-2025) series, comparing the 3B and 7B variants to assess scalability.
-Methods Accuracy
-Synthetic Real Overall
-ChartReasoner-3B
-Zero-Shot 46.41 42.90 45.90
-+CoT-SFT 51.02 45.45 50.22
-+GRPO-Zero 56.59 56.82 56.62
-+RFT 60.92 54.83 60.05
-ChartReasoner-7B
-Zero-Shot 60.54 45.17 58.34
-+CoT-SFT 63.59 51.70 61.88
-+GRPO-Zero 66.44 54.83 64.78
-+RFT 66.97 60.23 66.00
-Table 4: Ablation study showing the individual contribution of each strategy (CoT-SFT, GRPO-Zero, and RFT) to the final performance of the ChartReasoner models.
-The best and second-best scores in each column are highlighted using bold and underline formatting, respectively.
-Experimental Results
-Evaluation on CRBench
-The results in Table 1 underscore the significant challenge our CRBench presents to current models.
-The generally low scores across the board reveal a critical weakness: the widespread inability to reason over non-annotated charts.
-This difficulty stems from our benchmark’s emphasis on genuine visual reasoning—inferring values from axes and scales—rather than relying on annotation-based shortcuts common in other benchmarks.
-Furthermore, while most methods obtains comparable results on synthetic and real charts, the overall scores are still unsatisfying.
-To reveal the failure stems, we employed CoT prompting to compel models to externalize their reasoning process.
-The results in Table 1 and Table 2 show that most methods do not benefit from CoT prompting.
-It can be attributed to the lack of visual reasoning ability, as shown in Figure 3.
-Evaluation of ChartReasoner
-Our comprehensive evaluation confirms that ChartReasoner-7B decisively outperforms not only existing chart-specific models but also top-tier proprietary models like GPT-4o and Gemini-2.5-Flash, as shown in Table 2.
-This success stems from its ability to perform genuine visual reasoning, a skill that other models lack, as evidenced by their reliance on taking annotation shortcuts.
-This advanced capability is particularly evident in its handling of complex chart typologies.
-For instance, on challenging radar charts, where many models falter, ChartReasoner demonstrates a significant performance leap, proving it can interpret complex visual structures.
-Ablation Analysis of the Training Paradigm
-Our experimental results, summarized in Table 4, strongly validate the effectiveness of the RFT strategy for enhancing visual reasoning.
-As discussed in the previous section, simply prompting a model with a CoT format often leads to performance degradation, as it forces the model to articulate a reasoning process it does not possess.
-Our training paradigms were designed to address this fundamental issue.
-The result clearly indicates that while a naive CoT-SFT offers only a marginal improvement, it is the synergistic combination within the RFT model that yields the most significant performance gains, consistently outperforming all other configurations.
-Case Study of Visual Reasoning
+In contrast, our ChartVR-7B is able to produce a coherent and correct step-by-step reasoning process, leading to an accurate answer.
 In Figure 3, powerful models like GPT-4o and Qwen2.5-VL-7B either misinterpret the query or resort to factual hallucination.
-In contrast, our ChartReasoner, sculpted by the RFT strategy, demonstrates a flawless, step-by-step reasoning process.
-This case study provides a compelling visual proof of our quantitative findings: ChartReasoner enhanced by the RFT strategy is transformed from a system prone to errors and hallucination into a reliable and structured visual reasoner that moves beyond merely optimizing for the answer.
-Evaluation of ChartReasoner on the Public Benchmarks
-The robust generalization of ChartReasoner is confirmed by its strong performance across a suite of public benchmarks, as detailed in Table 3.
-This success stems from our model’s core visual reasoning capability, honed through our training strategy, which is highly transferable to other reasoning-intensive tasks.
-This approach contrasts with that of other chart-specific models relying on SFT.
-Although SFT can improve performance on specific tasks, it often fails to instill a generalizable reasoning faculty, limiting extensibility.
-In contrast, ChartReasoner, enhanced by the RFT strategy, is specifically designed to cultivate this fundamental reasoning, validating the superiority of our method.
-Conclusion
-In this paper, we investigate a compelling yet significant question: ”Do MLLMs really understand the charts?” By establishing the CRBench, we extensively evaluated open-source, close-source, and chart-specific MLLMs.
+In contrast, our ChartVR, sculpted by the RFT strategy, demonstrates a flawless, step-by-step reasoning process.
+This case study provides a compelling visual proof of our quantitative findings: ChartVR enhanced by the RFT strategy is transformed from a system prone to errors and hallucination into a reliable and structured visual reasoner that moves beyond merely optimizing for the answer.
+6 Conclusion
+In this paper, we investigate a compelling yet significant question: ”Do MLLMs really understand the charts?” By establishing the ChartVRBench, we extensively evaluated open-source, close-source, and chart-specific MLLMs.
 The results shows a significant degradation in the performance of these models, and, through chain-of-thought reasoning, revealed their inability to estimate numerical values through visual reasoning, similar to human behavior.
-To address this issue, we propose ChartReasoner, which enhances the visual reasoning ability of MLLMs via an RFT strategy.
+To address this issue, we propose ChartVR, which enhances the visual reasoning ability of MLLMs via an RFT strategy.
 This strategy first activates reasoning capabilities through SFT, and then generalizes reasoning abilities through RL.
-Extensive experiments conducted on the proposed CRBench and public chart reasoning datasets demonstrate the effectiveness of ChartReasoner.
+Extensive experiments conducted on the proposed ChartVRBench and public chart reasoning datasets demonstrate the effectiveness of ChartVR.
 This work paves the way for empowering MLLMs to really understand the charts in a human-like manner.
 Appendix
-  • Sec. A provides details of the proposed benchmark, CRBench, including the task definition, the data construction pipeline, and data sources, topics, and chart types.
-  • Sec. B describes the training strategy of the ChartReasoner, detailing the construction of the SFT and RL datasets.
-  • Sec. C presents further details on the inference.
-  • Sec. D includes other relevant details and extended results to supplement the main findings.
-Appendix A A. CRBench Benchmark Details
-Problem Definition
-The primary task addressed by our benchmark, CRBench, is numerical value estimation on non-annotated charts.
+  • Sec. A provides details of the proposed benchmark, ChartVRBench, including the problem definition, data sources, topics, and chart types.
+  • Sec. B describes the training strategy of the ChartVR, detailing the construction of the SFT and RL datasets.
+  • Sec. C presents further details on the evaluation and inference.
+  • Sec. D interpret the human performence in ChartVRBench.
+Appendix A ChartVRBench Benchmark Details
+A.1 Problem Definition
+The primary task addressed by our benchmark, ChartVRBench, is numerical value estimation on non-annotated charts.
 Formally, given a chart image $C$ and a query $Q$ that specifies a target data point within the chart, the goal is to produce a numerical answer $A$ that accurately estimates the value of that data point.
@@
 By designing a task that necessitates this entire reasoning chain, we directly evaluate a model’s ability to not just recognize a chart, but to truly understand its underlying quantitative information.
-Figure 4: The automated data generation pipeline for CRBench.
-Data Generation Pipeline
-The data generation pipeline for CRBench and training datasets, visualized in Figure 4, is an automated, multi-stage process designed to programmatically create a large and diverse set of non-annotated charts.
-The entire workflow is built upon generating and refining executable Python plotting scripts, ensuring that each data sample is verifiably correct.
+A.2 Data topics and Chart Examples
+Table 6: The 38 topics covered in the ChartVRBench dataset.
+Category Category
+Agriculture and Food Production Human Resources and Employee Management
+Architecture and Building Language and Communication
+Artificial Intelligence and Robotics Law and Legal Affairs
+Art and Design Literature and Writing
+Astronomy and Space Manufacturing and Production
+Biology and Life Sciences Marketing and Advertising
+Books and Publishing Mathematics and Statistics
+Business and Finance Music and Performance
+Computer Science and Information Technology Physics and Chemistry
+Education and Academics Real Estate and Housing Market
+Energy and Utilities Religion and Spirituality
+Environment and Sustainability Retail and E-commerce
+Fashion and Style Science and Engineering
+Film and Cinema Social Media and the Web
+Food and Beverage Industry Social Sciences and Humanities
+Futurism and Innovation Society and Community
+Government and Public Policy Sports and Entertainment
+Healthcare and Health Transportation and Logistics
+History and Culture Travel and Exploration
+To ensure the breadth relevance of ChartVRBench, the synthetic data generation process samples from a diverse set of 38 distinct topics, as shown in Table 6 guaranteeing that the charts cover a variety of contexts and narratives.
+Furthermore, to robustly evaluate a model’s visual reasoning capabilities across different graphical representations, ChartVRBench incorporates seven primary chart types.
+Figure 5 provides a representative example for each of these types, showcasing the visual diversity and complexity present in our benchmark.
+A.3 Real-World Chart Collection
+To ensure ChartVRBench reflects the challenges of real-world applications, we curated a substantial collection of charts from public online sources.
+This section details our three-stage process: data sourcing, a rigorous manual filtering protocol, and a hybrid human-AI pipeline for generating high-quality question-answer pairs.
+Figure 6 showcases several examples of the final curated real-world charts from our collection.
+A.3.1 Data Sourcing
+Statista.
+A significant portion of the real-world charts was sourced from Statista [^2] https://www.statista.com/, a leading global platform specializing in market and consumer data.
+Statista provides professional, data-driven visualizations for a wide array of industries, covering topics from economic indicators and market forecasts to technology trends and consumer behavior.
+Our World in Data.
+The second major source was Our World in Data [^3] https://ourworldindata.org/, a renowned scientific online publication based at the University of Oxford.
+Its mission is to make data and research on the world’s most significant challenges, such as global health, economic development, and environmental change, accessible and understandable through complex and data-rich visualizations.
+In addition to these two primary repositories, the collection was supplemented by charts from various other miscellaneous public reports and online publications.
+All data collection was conducted in strict adherence to the copyright policies, terms of service, and licensing agreements of each source to ensure full ethical compliance.
+A.3.2 Curation Protocol
+Once a large pool of charts was collected, each candidate chart underwent a meticulous, two-step curation process performed by our recruited team of human annotators.
+Step 1: Manual Filtering and Vetting.
+Each chart was manually vetted against stringent criteria for inclusion in ChartVRBench. A chart was accepted only if it satisfied all of the following conditions, otherwise it was discarded:
   1. 1.
-    Chart Generation by Large Language Model (LLM): The process begins with a curated seed pool of high-quality plotting scripts that serve as few-shot demonstrations.
-Using a Self-Instruct methodology, an LLM is then prompted to programmatically generate a vast and diverse set of new, non-annotated chart plotting scripts across 38 distinct topics.
+    High Visual Quality: The image must be of sufficient resolution, clear, and free of significant compression artifacts or other visual noise that could impede interpretation.
   2. 2.
-    Validation and Evolutionary Enhancement: Each generated script is rigorously validated via sandboxed execution to ensure it runs error-free.
-A key feature of the pipeline is an AI-driven repair loop, where the language model automatically debugs and corrects any code that fails execution.
-Subsequently, a portion of these validated scripts undergo an optional Evol-Instruct phase to further increase their visual and structural complexity.
+    Data Integrity: The chart must be coherent and visually consistent, with clearly defined axes, legends, and graphical elements.
   3. 3.
-    Final Curation: Only scripts that successfully pass the entire validation and enhancement pipeline are accepted into the final datasets.
-These validated scripts are then executed to generate the final, high-fidelity chart images used in our benchmark and datasets.
-Ground-Truth Q&A Generation and Verification.
-After generating the chart images, we create the corresponding ground-truth Q&A pairs.
-We first use a LLM to generate questions and answers based on the chart’s source code.
-Each generated Q&A pair then undergoes a rigorous, multi-tiered verification: it is first cross-checked for accuracy by another LLM instance, and subsequently undergoes a final manual review by human annotators to guarantee its unequivocal correctness.
-Data topics and Chart Examples
-To ensure the breadth relevance of CRBench, the data generation process samples from a diverse set of 38 distinct topics, span a wide range of domains including finance, healthcare, technology, and social sciences, guaranteeing that the charts cover a variety of contexts and narratives.
-Furthermore, to robustly evaluate a model’s visual reasoning capabilities across different graphical representations, CRBench incorporates seven primary chart types.
-Figure 6 provides a representative example for each of these types, showcasing the visual diversity and complexity present in our benchmark.
-Real-World Chart Collection
-To complement our synthetic dataset and ensure CRBench reflects the challenges of real-world applications, we curate a substantial collection of charts from public online sources.
-This process involves careful sourcing from key data providers, a rigorous manual filtering stage, and a hybrid human-AI approach to generate high-quality question-answer pairs.
-Statista
-A significant portion of the real-world charts is sourced from Statista [^2] https://www.statista.com/, a leading global platform specializing in market and consumer data.
-Statista provides professional, data-driven visualizations for a wide array of industries, covering topics from economic indicators and market forecasts to technology trends and consumer behavior.
-The charts on this platform are typically clean, well-structured, and designed for clear communication in business and academic contexts.
-The World in Data
-The second major source is Our World in Data [^3] https://ourworldindata.org/, a renowned scientific online publication based at the University of Oxford.
-Its mission is to make data and research on the world’s most significant challenges, such as global health, economic development, and environmental change, accessible and understandable.
-The visualizations from the World in Data are often more complex and data-rich, including long-term time-series line charts, multi-country scatter plots, and stacked area charts.
-In addition to these two primary repositories, the collection is supplemented by charts from various other miscellaneous public reports and online publications.
-All data collection is conducted in strict adherence to the copyright policies, terms of service, and licensing agreements of each source to ensure full ethical compliance.
-Manual Filtering and Vetting.
-A large pool of charts is initially collected.
-Each chart then undergoes a meticulous manual vetting process conducted by our team of annotators to ensure it met the stringent criteria for inclusion in CRBench.
-A chart is only accepted if it satisfied all of the following conditions:
+    Absence of Annotations: We exclusively select charts where numerical values are not explicitly printed on the graphical elements (e.g., no numbers on top of bars).
+This constraint is fundamental to our benchmark, as it forces a model to perform genuine visual reasoning rather than relying on OCR shortcuts.
+Step 2: Question-Answer Pair Generation.
+Once a chart was approved, we employed a two-stage, human-in-the-loop process to generate its corresponding question-answer pair:
   1. 1.
-    High Visual Quality: The image has to be of sufficient resolution, clear, and free of significant compression artifacts or other visual noise that could impede interpretation.
+    MLLM-based Candidate Generation: We first use a capable MLLM to generate an initial set of candidate question-answer pairs for each chart, prompting it to ask a specific numerical estimation question.
   2. 2.
-    Data Integrity: The chart has to be coherent and visually consistent, with clearly defined axes, legends, and graphical elements.
-  3. 3.
-    Absence of Annotations: We exclusively select charts where numerical values are not explicitly printed on the graphical elements (e.g., no numbers on top of bars or next to points on a line).
-This constraint is fundamental to our benchmark, as it forces a model to perform genuine visual reasoning rather than relying on OCR shortcuts.
-Charts that fails to meet any of these standards are discarded from the final dataset.
-Question-Answer Pair Generation.
-Once a real-world chart is approved, we employ a two-stage, human-in-the-loop process to generate its corresponding question-answer pair:
+    Human Verification and Refinement: Every MLLM-generated pair is then subjected to rigorous human review.
+Annotators verify the question’s clarity and relevance, and then carefully perform the visual estimation themselves to validate the answer’s accuracy.
+If necessary, they refine the question’s phrasing or correct the ground-truth answer to ensure the final Q&A pair is unambiguous and factually correct.
+Appendix B Details of Datasets and Training
+B.1 Data Construction for SFT
+The dataset used for the Supervised Fine-Tuning (SFT) “cold start” phase is meticulously constructed through a knowledge distillation process.
+The goal is to generate high-quality Chain-of-Thought (CoT) data that can effectively activate the reasoning paradigm of our base models.
+While we utilize the same underlying generation pipeline (rendering engine and topic distribution) to ensure domain consistency, the specific chart instances and question-answer pairs in the SFT set are distinct from those in the benchmark.
+Crucially, to maintain strict train-test separation, this SFT dataset is generated as a completely independent batch from the ChartVRBench evaluation set.
+This process leverages a powerful teacher model (Qwen2.5-VL-32B-Instruct) to generate reasoning traces for this training-specific corpus.
+The data construction pipeline involves several key stages to ensure the quality and validity of the final CoT samples:
   1. 1.
-    MLLM-based Candidate Generation: We use a capable Multimodal Large Language Model (MLLM) to generate an initial set of candidate question-answer pairs for each chart.
-The model is prompted to ask a specific numerical estimation question based on the visual information in the chart.
-  2. 2.
-    Human Verification and Refinement: Every MLLM-generated pair is subsequently subjected to rigorous human review.
-Annotators would first verify the question’s clarity and relevance.
-Then, they would carefully perform the visual estimation themselves to validate the accuracy of the answer.
-If necessary, they would refine the question’s phrasing or correct the ground-truth answer to ensure the final Q&A pair was unambiguous and factually correct.
-This careful, multi-stage curation process ensures that the real-world portion of CRBench is not only challenging and relevant but also contains reliable and accurate ground-truth data.
-Figure 7 showcases several examples of the final curated real-world charts from our collection, along with their corresponding question-answer pairs.
-Appendix B B. Details of Datasets and Training
-CoT Data Construction for Reasoning Activation
-The dataset used for the Supervised Fine-Tuning (SFT) ”cold start” phase is meticulously constructed through a knowledge distillation process.
-The goal is to generate high-quality Chain-of-Thought (CoT) data that can effectively activate the reasoning paradigm of our base models.
-This process leverages a powerful teacher model(In our experiment, we use the Qwen2.5-VL-32B-Instruct as the teacher mdoel) to generate reasoning traces for the question-answer pairs in our CRBench dataset.
-The data construction pipeline, implemented in the generation scripts, involves several key stages to ensure the quality and validity of the final CoT samples:
-  1. 1.
-    Prompting for Chain-of-Thought Generation: For each (C,Q,A) triplet from original data, we prompt the teacher model to generate a step-by-step reasoning process.
-Critically, the prompt(in Figure8) is designed to guide the model to derive the provided ground-truth answer, forcing it to articulate a logical pathway that leads to the correct result.
+    Prompting for Chain-of-Thought Generation: For each generated training instance, consisting of Python plotting code (C) and a question (Q), we employ the teacher LLM to generate a step-by-step reasoning process.
+The core of this process involves a carefully constructed textual prompt that integrates the Python code (C) and the question (Q).
+This prompt is designed to force the model to articulate a logical pathway from the code and question to the correct result.
 The model is required to structure its output using specific tags, separating the reasoning steps (<think>...</think>) from the final answer (<answer>...</answer>).
@@
     • Structural Check: The generated text is first parsed to ensure that both the reasoning and answer tags are present. Samples with missing tags are discarded.
-    • Answer Verification: The final answer extracted from the <answer> tag is programmatically compared against the ground-truth answer.
+    • Answer Verification: The final answer extracted from the <answer> tag is programmatically compared against the ground-truth answer derived from the code.
 We employ a robust evaluation function that checks for both exact string matches and numerical equivalence within a tolerance threshold to ensure correctness.
     • Leakage Detection: The generated reasoning trace within the <think> tags is scanned for any mention of the ”original answer.” This crucial step prevents the model from ”cheating” by simply copying the ground-truth answer into its reasoning, ensuring that the generated thought process is genuine.
-This meticulous, multi-stage distillation and verification process—with examples showcased in Figure 9 and Figure 10—ensures that the data used for the reasoning activation phase is both correct and logically sound, providing a strong foundation for the subsequent reinforcement learning stage.
-Dataset Construction for GRPO
-The dataset for the Generative Rejection-based Policy Optimization (GRPO) stage is meticulously curated to focus the training on the most informative examples.
-Instead of using a random subsample of data, we construct a specialized dataset of moderately difficult samples.
-These are question-answer pairs that the base model can sometimes answer correctly but frequently gets wrong, representing the frontier of its reasoning capabilities and thus offering the richest signal for reinforcement learning.
-The curation process inspired by the principles of rejection sampling and active learning, involves a multi-round, varied-prompting inference pipeline designed to probe the model’s knowledge boundaries.
-We first run inference on the entire CRBench dataset multiple times using the same base model.
-To elicit diverse responses, we systematically vary the prompts between rounds (e.g., alternating between direct answering and Chain-of-Thought prompts) and increase the model’s sampling temperature to 0.9.
-The correctness of every generated response is logged.
-After the evaluation is complete, we filter this log to isolate the key samples: we select only those question-answer pairs that the model answered correctly in at least one round but incorrectly in at least one other round.
-The rationale behind this selective filtering strategy is to maximize the efficiency and impact of the subsequent RL training.
-By focusing on these ”boundary” cases, we avoid wasting computational resources on problems the model has already mastered (always correct) or those that are currently far beyond its reach (always incorrect).
-Training on these ”boundary” samples forces the policy to learn to distinguish between successful and flawed reasoning pathways for the exact same problem.
-This targeted approach ensures that the GRPO stage is dedicated to resolving ambiguity and reinforcing robust reasoning where it is most needed, leading to more significant and generalizable improvements in the model’s core reasoning abilities.
-Detailed Formulation of the Accuracy Reward
-A core component of the GRPO framework is the Accuracy Reward ($R_{\text{acc}}$), which is designed to provide a dense, fine-grained signal for optimizing the model’s numerical estimation capabilities.
+B.2 RL Algorithm Selection
+We selected GRPO to fine-tune our multimodal model for enhanced chart visual reasoning, primarily due to its superior efficiency and its alignment with our reward structure.
+Compared to Proximal Policy Optimization (PPO), GRPO significantly reduces computational and memory overhead.
+GRPO eliminates the need for a separate value model—which is typically as large as the policy model—by estimating the baseline directly from the scores of multiple sampled outputs.
+This efficiency is critical given the large scale of our models (e.g., 7B parameters), making GRPO a practical solution under limited hardware resources.
+Furthermore, while Direct Preference Optimization (DPO) offers an efficient alternative to traditional RLHF, it is fundamentally designed for binary preference datasets (i.e., chosen vs.
+rejected responses).
+Our task, however, benefits from a more granular, continuous reward signal that reflects the degree of correctness in quantitative analysis.
+GRPO is adept at directly optimizing for such programmatic, scalar rewards, allowing the model to learn from fine-grained feedback.
+This makes it better suited for improving the visual reasoning in chart than a preference-based method like DPO.
+B.3 Dataset Construction for GRPO
+The curation process, inspired by the principles of rejection sampling and active learning, involves a multi-round, varied-prompting inference pipeline designed to probe the model’s knowledge boundaries.
+The goal of this pipeline is to construct a specialized, high-signal dataset by isolating ambiguous cases that the SFT-tuned model can sometimes solve but not consistently.
+This strategy focuses the training process on the most informative examples where the model is most uncertain, rather than wasting computational resources on problems that are already mastered (always correct) or are currently too difficult (always incorrect).
+Our pipeline involves the following systematic steps:
+  1. 1.
+    Initial Dataset Curation: We begin by constructing an initial, high-quality dataset for GRPO.
+This dataset is synthesized following the method introduced earlier.
+We selected only those instances that achieved a score of either 4 or 5, ensuring a strong baseline of correct.
+  2. 2.
+    Multi-Round Inference: Initially, we run inference multiple times on the initial GRPO dataset using our base model.
+To elicit a wide range of reasoning pathways and outcomes for each problem, we set the sampling temperature to 1.0 for each run.
+  3. 3.
+    Filtering for ”Stochastic Correctness”: The correctness of every generated response is logged.
+After all rounds are complete, we filter this log to isolate the target samples.
+We select only those question-answer pairs that the model answered correctly in at least one round but incorrectly in at least one other round.
+The rationale behind this selective filtering is to force the policy to learn to distinguish between successful and flawed reasoning pathways for the exact same problem.
+Training on these ”boundary” cases ensures that the GRPO stage is dedicated to resolving ambiguity and reinforcing robust reasoning where it is most needed, leading to more significant and generalizable improvements in the model’s core abilities.
+B.4 Detailed Formulation of the Accuracy Reward
+A core component of the GRPO framework is the Continuous Accuracy Reward ($R_{\text{acc}}$), which is designed to provide a dense, fine-grained signal for optimizing the model’s numerical estimation capabilities.
 A simple binary reward (correct/incorrect) is often too sparse for reinforcement learning, as it fails to differentiate between a near-miss and a completely wrong answer.
 To overcome this, we designed a continuous reward function that recognizes and rewards ”nearly correct” answers, thereby creating a smoother optimization landscape.
-The formulation of this reward is visualized in Figure 5.
-Figure 5: Visualization of our quadratic accuracy reward function ($R_{\text{acc}}$) with a tolerance threshold of $\varepsilon=0.02$.
-The reward smoothly decays from 1 to 0, with a gentle slope for small errors and an accelerating penalty for larger deviations.
-For a predicted numerical answer $A_{\text{pred}}$ and a non-zero ground-truth answer $A_{\text{gt}}$, we first calculate the relative error:
-$d_{\text{rel}}=\frac{\lvert A_{\text{pred}}-A_{\text{gt}}\rvert}{\lvert A_{\text{gt}}\rvert}$ (6)
-Based on a tolerance threshold of $\varepsilon=0.02$, we define the reward using a piecewise quadratic function that smoothly decays from a maximum of 1 to 0:
-$R_{\text{acc}}(d_{\text{rel}})=\begin{cases}\left(1-\frac{d_{\text{rel}}}{\varepsilon}\right)^{2}&\text{if }d_{\text{rel}}<\varepsilon\\ 0&\text{if }d_{\text{rel}}\geq\varepsilon\end{cases}$ (7)
 Our accuracy reward function provides a dense, informative, and well-behaved signal that is ideally suited for guiding our reinforcement learning process towards generating highly accurate and reliable numerical estimations.
-Appendix C C. Inference Details
-This section provides additional details regarding our inference and evaluation setup to ensure full reproducibility.
-Inference Hyperparameters and Framework
+B.5 More Training Details
+Our training process begins with the Qwen2.5-VL-7B-Instruct model as the foundation.
+Using the Swift framework, we perform SFT for 2 epochs on our 4.2k instruction-following dataset.
+In this stage we freeze the vision tower and the multimodal aligner while exclusively tuning the LLM’s parameters.
+We set the learning rate to 1e-5 with a warm-up ratio of 0.05 and use an effective batch size of 256.
+The SFT process is conducted on 8 NVIDIA H800 GPUs, utilizing bfloat16 precision and the DeepSpeed ZeRO-3 optimization strategy.
+For the GRPO stage, we initialize the model with the checkpoint from the SFT phase and employ the GRPO algorithm on our 3.4k preference dataset.
+In this phase, we continue to freeze the vision tower but expand the scope of fine-tuning to include both the LLM and the multimodal aligner.
+The learning rate is reduced to 1e-6, again with a 0.05 warm-up ratio.
+For the rollout process, we use a generation batch size of 32 to create 4 completions per sample with a temperature of 1.0; the training itself uses an effective batch size of 64.
+The reward function is the composite of the Format and Continuous Accuracy rewards.
+The hardware and optimization setup remains consistent, utilizing 8 NVIDIA H800 GPUs with bfloat16 precision and DeepSpeed ZeRO-3.
+Appendix C Evaluation and Inference Details
+This section outlines the precise methodologies and inference settings used to evaluate all models and benchmarks, ensuring full reproducibility and fairness.
+Our protocols were designed by strictly adhering to the author-recommended settings and official evaluation scripts where available.
+C.1 General Inference Settings
 All experiments reported in this paper were conducted using the default hyperparameters of each respective model, with no model-specific tuning performed at inference time.
-This ensures a fair and consistent comparison across all models.
-To guarantee reproducibility, the random seed for all generation processes was fixed to 42.
+To ensure reproducibility, the random seed for all generation processes was fixed to 42, and the sampling temperature was set to 1.0.
 Our inference pipeline is built upon the vLLM framework, which provides efficient, high-throughput serving for Large Language Models.
-Prompting Strategies for Evaluation
-To thoroughly investigate the models’ reasoning capabilities and behaviors, we employed three distinct prompting strategies during evaluation, designed to control the externalization of the model’s thought process.
-The first is Direct Answering, a baseline prompt that instructs the model to provide only the final answer.
-The second, Optional Chain-of-Thought (Optional-CoT), encourages but does not require the model to generate a reasoning trace, allowing us to observe its natural tendency to reason.
-The third, Forced Chain-of-Thought (Forced-CoT), strictly requires a step-by-step reasoning process before the final answer, compelling the model to articulate its logic.
-As illustrated in Figure 14 and Figure 15, which showcases two comparative examples, these varied strategies allow us to gain deeper insights into not just what the models answer, but how their performance is affected by the explicit demand for reasoning.
-Appendix D D. Evaluation Protocols for Specialized Models and Benchmarks
-This section outlines the precise met hodologies used to evaluate specialized, chart-specific models and to benchmark our ChartReasoner on existing public datasets.
-Our protocols were designed to ensure maximum fairness and comparability by strictly adhering to the author-recommended settings and official evaluation scripts for each model and benchmark.
-Evaluation of Chart-Specific Models on CRBench
-Evaluating the performance of existing chart-specific models on our proposed CRBench required a careful approach to ensure a fair comparison.
-These models, including ChartGemma, TinyChart, ChartInstruct, ChartVLM, and ChartLlama, are often fine-tuned on specific data formats and are highly optimized for a particular style of prompting.
-For this reason, we did not use our generalized prompting templates (as shown in Figures 12 and 13) for these models.
-Instead, to elicit the best possible performance from each model and establish the strongest baseline, we followed this procedure:
+C.2 Evaluation on ChartVRBench
+Our evaluation on the proposed ChartVRBench employed different prompting strategies depending on the model type to ensure a fair and rigorous assessment.
+General MLLMs.
+To verify the visual reasoning capabilities of general-purpose models, we employed a structured Chain-of-Thought (CoT) prompt, shown in Figure 10.
+This prompt compels the model to first articulate its reasoning—by analyzing axes, data points, and context—before providing a final answer.
+The prompt enforces a strict separation between the step-by-step logic (output in ‘¡think¿‘ tags) and the concise final output (in an ‘¡answer¿‘ tag).
+This approach allows us to pinpoint the exact stage where a model’s logic succeeds or fails, moving our analysis beyond simple accuracy metrics.
+Chart-Specific Models.
+In contrast, for models already fine-tuned on specific chart-related data formats (including ChartGemma, TinyChart, ChartInstruct, ChartVLM, and ChartLlama), we did not use our generalized CoT prompt.
+To elicit their best possible performance and establish the strongest baseline, we followed the official author-recommended procedures:
   1. 1.
-    We cloned the official public repository associated with each chart-specific model.
+    We cloned the official public repository for each model.
   2. 2.
@@
   3. 3.
-    We fed the images and questions from our CRBench test set into these scripts.
-This methodology ensures that the results reported in the main body of the manuscript represent each model performing under its own optimal, author-intended conditions.
-This prevents any potential performance degradation that might arise from imposing a foreign prompt structure and guarantees that we are comparing ChartReasoner against the most capable version of each specialized model.
-Evaluation of ChartReasoner on Public Benchmarks
-To validate the generalization of ChartReasoner’s enhanced reasoning capabilities, we evaluated it against several standard public benchmarks.
-The following subsections detail the exact procedure used for each.
-CharXiv
-For the CharXiv benchmark, we utilized the official code and evaluation scripts available in their repository.
-Our evaluation process involved adapting their provided inference pipeline.
-We integrated our locally-deployed ChartReasoner (served via vLLM) as the model backend, replacing the default model call.
-All other components, including data loading, pre-processing, and the final scoring scripts for both the reasoning (R) and descriptive (D) tasks, were kept identical to the original setup.
-ChartBench
-Our evaluation on ChartBench followed a similar protocol.
-We used the complete evaluation pipeline provided in the official ChartBench repository.
-Predictions for each test instance were generated by passing the questions to our ChartReasoner vLLM endpoint.
-The resulting outputs were then fed directly into the official scoring script, which computes the overall accuracy based on their defined metrics and question categories.
-ChartQAPro
-The official repository for ChartQAPro provides a standalone evaluation script but does not include a full inference pipeline.
-Therefore, to benchmark ChartReasoner, we implemented a two-step process that carefully reconstructs the intended evaluation conditions: 1.
-Prediction Generation: We developed a custom script to generate predictions for the entire ChartQAPro test set.
-For each question, this script formatted a query for ChartReasoner that precisely replicated the direct prompt template described in the official ChartQAPro paper.
-2.
-Scoring: The complete set of generated predictions was saved to a file.
-This file was then used as input for the official evaluation script provided by the ChartQAPro authors.
-The script automatically compares our model’s predictions against the ground-truth annotations to compute the final accuracy score.
-Figure 6: An overview of sample question-answer pairs for various synthetic chart types within the CRBench dataset.
-Figure 7: An overview of sample question-answer pairs for various real chart types within the CRBench dataset.
+    We fed the images and questions from our ChartVRBench test set directly into these scripts.
+This methodology ensures that we are comparing our model against the most capable version of each specialized baseline.
+C.3 Evaluation of ChartVR on Public Benchmarks
+To validate the generalization of ChartVR’s enhanced reasoning capabilities, we evaluated it against several standard public benchmarks, following the official protocol for each.
+CharXiv.
+We utilized the official code and evaluation scripts from the CharXiv repository.
+We integrated our locally-deployed ChartVR as the model backend into their inference pipeline, keeping all other components (data loading, pre-processing, and scoring scripts) identical to the original setup.
+ChartBench.
+Our evaluation followed a similar protocol using the complete pipeline from the official ChartBench repository.
+We generated predictions with our ChartVR model and fed the outputs directly into the official scoring script.
+ChartQAPro.
+As the official repository provides a standalone evaluation script but not a full inference pipeline, we implemented a two-step process.
+First, we developed a script to generate predictions for the test set using a prompt that precisely replicated the template described in the ChartQAPro paper.
+Second, the resulting file of predictions was used as input for the official evaluation script to compute the final accuracy score.
+Appendix D Interpretation of Human Performance
+It is important to note that human accuracy on this task is not 100%.
+This is primarily due to inherent tendencies in human visual estimation, for instance, individuals often gravitate towards estimating with round or integer values that appear close to the correct answer, rather than performing precise interpolation.
+Our analysis indicates that a 2% relative error tolerance is a reasonable threshold to account for these natural human inaccuracies.
+Furthermore, performance varies significantly across chart types.
+For area charts, accuracy sees a substantial decline, because many questions require calculating the difference between the upper and lower boundaries of a shaded region, a task made considerably more difficult by the common absence of horizontal gridlines as visual aids.
+Similarly, for complex combo charts, lower performance can be attributed to cognitive factors, such as misinterpretation of the prompt or misunderstanding the intricate relationships between different chart components.
+Figure 4: An overview of sample question-answer pairs for various synthetic chart types within the ChartVRBench dataset.
+Figure 5: An overview of sample question-answer pairs for various complex synthetic chart types within the ChartVRBench dataset.
+Figure 6: An overview of sample question-answer pairs for various real chart types within the ChartVRBench dataset.
 Figure 7: The Prompt in Distilling the CoT Data
@@
 Figure 9: Example 2 of CoT Dataset
-Figure 11: Direct Prompting in CRBench Evaluation
-Figure 12: Prompting for Optional Output of Reasoning Steps in CRBench Evaluation
-Figure 13: Prompting for Forced Output of Reasoning Steps in CRBench Evaluation
-Figure 14: Example 1 from CRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartReasoner-7B.
-Figure 15: Example 2 from CRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartReasoner-7B.
+Figure 10: CoT Prompting Reasoning Steps in ChartVRBench Evaluation
+Figure 11: Example 1 from ChartVRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartVR-7B.
+Figure 12: Example 2 from ChartVRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartVR-7B.